A Complete Bibliography of Publications in Technometrics for the decade 2020–2029

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

13 October 2023 Version 1.08

Title word cross-reference

$102.12 [Els20]. $107.78 [Wlu21]. $118.00 [Mal23a]. $119.35 [Mye21]. $119.99 [Lip20d]. $120.00 [Lip23f]. $123.83 [Ahm21d]. $129.00 [Lip20h, Lip21l]. $130.00 [Lip21m, Lip22j, Mal23c]. $135 [Nel21]. $135.00 [Lip21b]. $136.46 [Sau23]. $136.93 [Ahm21c]. $140.00 [Che22, Lip21h, Ndo24]. $152.00 [Lip21g]. $160.00 [Lip20a, Lip21c]. $174.95 [VL21b]. $179.99 [Ahm20a]. $185.15 [Zel22]. $19.99 [ZGAR24]. $200.00 [Lip23c]. $23.96 [Lip23d]. $23.99 [Kha23]. $24.95 [Fir23]. $25.00 [Mal22a]. $28.00 [Lip22b]. 2^{5} [Lip21a]. 3 [OKF22, SA23b]. $31.16 [Lip23g]. $34.00 [Lip23a]. $37.99 [Lip20c]. $38.00 [Lip22a]. $39.99 [Lip23b]. $46.36 [Lip21g, Lip21h, Wlu21]. $49.99 [Lip22h]. $50 [Ano21d]. $50.78 [Mal20]. $52.49 [Mal23b]. $53.59 [MD24]. $54.95 [Lip22n]. $55.96 [Lip22i]. $59.96 [Lip21e, Lip21d]. $59.98 [Ahm20d]. $63.96 [Lip22c, Lip22g]. $64.99 [Mal23d]. $66.95 [Lip22m]. $68 [Mal23e]. $69.99 [Lip20f, Lip24a]. $74.99 [Ano21b, Lip20e]. $75.95 [Lip21f]. $76.64 [DM20]. $79.95 [Lip20i, Lip22k, Oli20, Wlu23]. $79.96 [Lip22e, Lip22l, Lip22f]. $83.99 [Ahm21e]. $87.20 [Cha22]. $87.22 [Lu20]. $88 [Liu24]. $88.00 [Asz23]. $89.00 [Lip20j]. $89.95 [Lip22d]. $89.99 [Lip24b]. $93.18 [Ano21a]. $94.94 [Mal21]. $95.41 [Mar20]. $99.00 [Mal22b]. $99.95 [Ano21e, AL22, Lip20b, Lip20g, Lip21j, Lip21k, Oli22, Sea20, Lip21i]. $99.99 [Ahm20b]. _{2} [HCL23]. A [SAMJ23]. D [NPM20, SAMJ23]. k [LZC21]. t [PC20]. Var(s) [WSSE21]. φ_{p} [VE21].

[Ano21e].

-Efficient [NPM20]. -Layer [LZC21]. -Optimal [SAMJ23].

//doi.org/10.1201/9781003187158 [DSB23].

0 [APF24, Lip21g, Lip23g, Ndo24, Sea20].

1 [Ahm20a, Ahm21c, Lip20i, Lip21e, Lip22f, Lip24a, Lip22a]. 112.00 [Ano21c]. 13 [Mye21]. 13-4822-3742-9 [Mye21]. 1st [APF24, SA23c].

2 [Hus23, Lip20h, Lip22a, Mal22a, Mal23a, Mal23e, Wlu23, Lip21e]. 2018 [Ano20a, Ano20b]. 2020 [Che22]. 2021 [Ano22a]. 2022 [Ano22b]. 2nd [Ahm20b, Ahm20c, Ahm21b, AL22, Fir23, Lip21c, Lip22i, Lip22j, Lip23e, Lu20, Oli20, Sau21].

3 [Ahm20c, DM20, Lip20d, Lip20e, Lip21f, Lip21m, Lip23d, Mal22b, Mal20, Sa’23a]. 3rd [Lip22n, Mal22a].

4 [Ahm20b, Ahm21e, Lip22j, Lip23f, Mal23c, Zel22]. 4.0 [Pra22].

5 [Fir23, GC23, Lip20c, Lip22l, Lip22i, Lip23a, Oli22]. 5th [Mal22b].

6 [Ahm20d, Els20, Lip20b, Lip20g, Lip21b, Lip23c, Mal22c, MD24]. 6th [Ano21b, Lip20g].

7 [Ahm21d, Gho20, Kha23, Lip21l, Lip22d, Lip22k, Lip24b, Lu20, Rah23, ZGAR24].

8 [AL22, Lip20a, Lip21d, Lip21j, Lip23b, Liu21, Mal23b, Mal23d, VL21b, Wlu21].

9 [Lip20f, Lip20j, Lip21c, Lip22b, Lip22e, Mal21, Mye21, Nel21, Oli20]. 978 [Ahm20a, Ahm20b, Ahm20c, Ahm20d, Ahm21c, Ahm21d, Ahm21e, AL22, APF24, DM20, Els20, Fir23, GC23, Gho20, Hus23, Kha23, Lip20a, Lip20b, Lip20d, Lip20c, Lip20e, Lip20h, Lip20f, Lip20g, Lip20i, Lip20j, Lip21b, Lip21c, Lip21g, Lip21f, Lip21e, Lip21d, Lip21j, Lip21l, Lip21m, Lip22d, Lip22a, Lip22b, Lip22e, Lip22l, Lip22f, Lip22i, Lip22k, Lip22j, Lip23a, Lip23b, Lip23f, Lip23c, Lip23d, Lip23g, Lip24a, Lip24b, Liu21, Lu20, Mal22a, Mal22b, Mal22c, Mal23a, Mal23b, Mal23c, Mal23d, Mal23e, MD24, Mal20, Mal21, Ndo24, Nel21, Oli20, Oli22, Rah23, Sa’23a, Sea20, VL21b, Wlu21, Wlu23, Zel22, ZGAR24]. 978-0-367-07611-5 [Lip22i]. 978-0-367-13559-1 [Lip22f]. 978-0-367-13991-9 [Mal21]. 978-0-367-18061-4 [Mal23c]. 978-0-367-19149-8 [AL22]. 978-0-367-22104-1 [Lip21e]. 978-0-367-22118-8 [Lip21d]. 978-0-367-22120-3 [Lip21f]. 978-0-367-25538-1 [Lip21i]. 978-0-367-34014-8 [Ano21c]. 978-0-367-35797-9 [Oli20]. 978-0-367-35798-6 [Lip20b]. 978-0-367-36292-8 [Wlu21]. 978-0-367-40820-6 [Lip21h]. 978-0-367-40878-7 [Lip22d]. 978-0-367-40982-1 [Lip20i]. 978-0-367-42452-7 [Lip22k]. 978-0-367-46094-5 [Lip22l]. 978-0-367-47230-6 [Liu24]. 978-0-367-47407-2 [Liu24]. 978-0-367-53891-0 [Lip22m]. 978-0-367-54471-3 [Lip23d]. 978-0-367-57360-7 [Lip22n, Ahm21d]. 978-0-367-70815-3 [Mal22b]. 978-0-367-74090-0 [Lip23g]. 978-0-367-75901-8 [Mal23b]. 978-0-367-76344-2 [Hus23]. 978-0-367-81889-0 [Lip21k]. 978-0-429-34051-2 [Wlu23]. 978-0-471-68717-7 [Lu20]. 978-0-691-17762-5 [Fir23]. 978-0-691-18237-7 [Rah23]. 978-0-8153-7468-8 [Lip20a]. 978-0-982-91746-6 [Ano21d]. 978-0-9829174-7-3 [Ahm20c]. 978-03-6734-187-9 [Lip21c]. 978-1-003-03534-3 [Liu24]. 978-1-032-00910-0 [APF24]. 978-1-032-03404-1 [DSB23]. 978-1-032-04174-2 [Lip22g]. 978-1-032-06536-6 [Asz23]. 978-1-032-06704-9 [Lip22e]. 978-1-032-12578-7 [Kha23]. 978-1-032-15373-5 [Oli22]. 978-1-032-16379-6 [MD24]. 978-1-032-20821-3 [SA23c]. 978-1-032-24407-5 [GC23]. 978-1-0716-2986-4 [HL23]. 978-1-108-82341-8 [Lip23b]. 978-1-108-95998-8 [Mal23d]. 978-1-118-11545-9 [Nel21]. 978-1-118-31103-5 [Che22]. 978-1-118-38798-6 [Lip21b]. 978-1-118-59448-3 [Nel21]. 978-1-118-59459-9 [Nel21]. 978-1-119-42231-0 [Ndo24]. 978-1-119-71918-2 [Mal22a]. 978-1-138-05400-4 [Ano21a]. 978-1-138-10640-6 [Lip23c]. 978-1-138-39392-9 [Ano21e]. 978-1-138-48395-8 [Lip21j]. 978-1-138-49594-4 [Lip22j]. 978-1-138-59052-6 [Els20]. 978-1-138-60733-0 [Lip21g]. 978-1-138-70222-6 [Lip20g]. 978-1-138-71533-2 [Lip22c]. 978-1-1380-3232-3 [Lip21m]. 978-1-352010-25-1 [Ano21b]. 978-1-4398-7651-0 [Sea20]. 978-1-482-24274-4 [Lip23f]. 978-1-49-877321-8 [VL21b]. 978-1-4939-8853-2 [Cha22]. 978-1-4987-1462-4 [Zel22]. 978-1-68358-466-7 [ZGAR24]. 978-3-030-02461-1 [Lip24a]. 978-3-030-33415-4 [Ahm21e]. 978-3-030-45733-4 [Lip22h]. 978-3-030-47204-7 [Lip24b].

978-3-030-55799-7 [Lip21l]. 978-3-030-59376-6 [Mal22c]. 978-3-319-71391-5 [Lip20c]. 978-3-319-78965-1 [Ahm20a]. 978-3-319-93175-3 [Lip20e]. 978-3-319-93631-4 [Ahm20b]. 978-3-319-94152-3 [Mal20]. 978-3-319-94988-8 [Liu21]. 978-3-319-95663-3 [DM20]. 978-3-319-99388-1 [Ahm21c]. 978-319-78827-2 [Lip20h]. 978-319-91739-9 [Lip20f]. 978-319-92581-3 [Lip20d]. 978-319-95419-6 [Ahm20d]. 978-319-97063-9 [Lip20j]. 978-981-12-2155-2 [Mal23e]. 978-981-12-2650-2 [Lip22a]. 978-981-12-2651-9 [Lip22b]. 978-981-12-3710-2 [Mal23a]. 978-981-12-4783-5 [Lip23a]. 978-981-13-1549-7 [Gho20]. 978-981-13-2145-0 [Mar20]. 978-981-13-2146-7 [Mar20]. 978-981-16-8615-3 [Sa’23a]. 978-981-19-1043-2 [Liu23]. 9781003187158 [DSB23].

A/B [ZK22]. Abdelmonem [Lip20g]. Abdominal [SBWK22]. Abu [Ndo24]. Abu-Rub [Ndo24]. Accelerated [KTF21, WM21, WV23]. Accounting [MCD^{+}22]. Accuracy [LJK20]. Achieve [VE21]. Achieving [Ano21a]. Acquisition [SWL^{+}22]. Active [CKL21, SGH23]. Adam [Mal23d]. Adams [Lip21i]. Adaptation [SA23b]. Adaptive [AGC^{+}21, CJXD21, GLP22, LC23, WXJT21, YQ21, YXW23, ZWX23, ZHMY22]. Addition [LC23, Tsa23, YSX21]. Additive [GLLL21, STC^{+}22, SKD20, SWL^{+}22]. Additivity [LJ21]. Advanced [Ano21b, SA23c, Ano21b, Lip21b, Sau21]. Advances [Lip20d]. Advection [LY23]. Advection-Diffusion [LY23]. Adventures [Lip22a, Lip22b]. Afifi [Lip20g]. After [Zou21]. AG [Lip21l, Lip22h, Lip24a, Lip24b, Liu21]. Age [Lip23b]. Aggregate [WY21]. Agresti [Sau23]. Ahmed [Hus23]. Ahn [Lip22l]. AI [DSB23]. Aided [PTD22]. Aircraft [LLK^{+}23]. Aircraft-UAV [LLK^{+}23]. Alan [Lip20c, Sau23]. Albert [Lip20i]. Alexander [Lip21h]. Alfred [Mal23a]. Algaba [Lip20a]. Algorithm [LC23]. Algorithms [Lip20b, Lip23b]. Aliased [NG20]. Alignment [QSK21]. Almela [Lip20d]. Almela-Sanchez [Lip20d]. Almost [Lip24a]. Alvo [Mal20]. Analysis [AX20, Ahm20a, Ahm20d, Ahm21d, ECNA20, GSSI22, HSZK20, JLM^{+}21, KK21, LL21, LZC21, LCX23, Lip20b, Lip20d, Lip22e, Lip22f, Lip22j, Lip22n, LP20, Mal23b, PC20, SKD20, TCXM23, WV21, WZ22, WWYC24, YSX21, ZPS23, ZMXX23, Lip20g]. Analytic [Lip21m, Ahm21b]. Analytics [DM20, Lip22g]. Analyzing [BA22, Lip22c]. Anandarajan [DM20]. Ancient [Mal23e]. Andrea [Lip20h]. Andrés [Lip21k]. Andries [Lip22k]. Aneurysm [SBWK22]. Ann [VL21b]. Announcement [Ano20b, Ano22c, Ano22b]. Anomalies [HTLS21]. Anomaly [GMLB22, LMN^{+}22]. Anthony [Mal23c]. Antonelli [Liu21]. Antony [Mal22b]. Aortic [CMJZ21, SBWK22]. Application [AX20, AS21, CKJD23, CJXD21, KK21, LJK20, LCL23a, LDY23, MCD^{+}22, STC^{+}22, SA23b, WZZ23, YPLS20, ZSPT21]. Applications [CMJZ21, Gho20, Lip21b, Lip21d, Lip22l, Liu23, LLK^{+}23, PTD23, SWL^{+}22, VL21a, Ahm20d, Lip20d, Lip22j]. Applied [Pou21, SBWK22, Lip22d]. Approach [FYQ^{+}21, GYPS21, GSSI22, Gho21b, GYZ23, LH22, LP23, LT22, Liu21, Mal20, OKF22, PL21, WDZS22, Xio21, ZTCA21, ZdC21, Mye21, SA23c]. Approaches [Ahm20d, Lip20h]. Arboretti [Lip20f]. Areas [Lip20c]. Arjuna [Lip20j]. Ark [Lip22k]. Armstrong [Lip22c]. Arnold [Liu23]. Arrays [Tsa23]. Art [Lip22i]. Artificial [MMO24, Sa’23a, Wlu21]. Artioli [Mal23e]. Ashis [Liu23]. Asis [Mar20]. Assembly [DCH^{+}22]. Assisted [CKL21]. Associate [Ano24]. Association [Gho21b]. Assurance [WF21]. Atomic [MCD^{+}22]. Attributes [LP20, Lip22k]. Augmentation [ZWX23]. Aurore [Lip23c]. Auto [CMAKA23]. Auto-Regressive [CMAKA23]. Autoregressive [YSM23]. Award [Ano20b, Ano22b]. Awari [MD24].

B [DSB23, Lip23f, Lip23c, Lip23g, Lip21i, Lip21j, Lip21k, Lip22d, Lip22e, Lip22h, Lip22k, Lip24b, Liu23, Liu24, ZK22]. B/W [DSB23, Lip23f, Lip23c, Lip23g, Lip21i, Lip21j, Lip21k, Lip22d, Lip22e, Lip22h, Lip22k, Lip24b, Liu23, Liu24]. Baccarat [Lip23g]. Backseat [Liu24]. Bahadir [Hus23]. Bailer [Oli20]. Bakker [Lip22c]. Bandit [ZM23]. Banerjee [Lip22j]. Barry [Liu23]. BART [HSSP22]. Base [WWYC24]. Baseball [Ahm21b]. Based [AX20, DSW23, LT22, WV21, WDZS22, WZL22, ZWX23, ZH21, ZPS23, MF21, WT23, XSL^{+}23, ZSD24]. Basic [Mal22b]. Basketball [Ahm21b]. Bass [Mal22a]. Bathke [Lip20f]. Baumer [AL22]. Bayes [Lip21o]. Bayesian [Lip20c, Mal21, Mye21, ABRA20, CHH^{+}23, CKJD23, GR21, Gho21c, GG21, GYZ23, GG22, HPR^{+}21, JYYJ22, KK21, LJK20, LK21, LT22, MF21, QBN21, RD24, WM21, XKSK21, XSL^{+}23]. Bayhan [Ndo24]. Begovic [Ndo24]. Behavior [Lip22e]. Behind [ZGAR24]. Being [Ano21e]. Benjamin [AL22]. Bernd [Mal23a]. Better [Ahm21d, Lip22n]. between [YXW23]. Bias [GAM23, WXJT21]. Bicentennial [Liu23]. Biecek [Lip22f]. Big [Ahm21d, Ano21e, GG22, Lip22n, Lip23g, LP20, ZWX23, Lip21c]. Bikas [Mar20]. Binary [LJ22]. Bioinformatics [Ahm21c]. Biology [SA23c]. Biomedical [Ahm21e]. Biondi [Ahm20a]. Biondi-Zoccai [Ahm20a]. Biostatistics [Ahm21c, Wlu23]. Bivariate [AS21]. Blaabjerg [Ndo24]. Blackjack [Lip23g]. Blocked [MGG20]. Blocking [SYH21]. Blockmodel [PHRD22]. Boca [Ahm21d, Ano21a, Ano21e, AL22, Asz23, DSB23, Els20, GC23, Lip20a, Lip20b, Lip20g, Lip20i, Lip21c, Lip21g, Lip21f, Lip21e, Lip21d, Lip21h, Lip21i, Lip21j, Lip21k, Lip21m, Lip22c, Lip22d, Lip22e, Lip22l, Lip22f, Lip22g, Lip22i, Lip22k, Lip22m, Lip22n, Lip23f, Lip23c, Lip23d, Lip23g, MD24, Mal21, Mye21, Oli20, Oli22, Sau23, Sea20, SA23c, VL21b, Wlu21, Wlu23, Zel22]. Boka [Lip22j]. Bollman [Lip23g]. Bonnini [Lip20f]. Book [Ahm20a, Ahm20b, Ahm20c, Ahm20d, Ahm21b, Ahm21a, Ahm21c, Ahm21d, Ahm21e, Ano21a, Ano21b, Ano21c, Ano21d, Ano21e, AL22, APF24, Asz23, Cha22, Che21, Che22, DM20, DSB23, Els20, Fir23, GC23, Gho20, HL23, Hus23, Kha23, Lip20a, Lip20b, Lip20d, Lip20c, Lip20e, Lip20h, Lip20f, Lip20g, Lip20i, Lip20j, Lip21b, Lip21c, Lip21g, Lip21f, Lip21e, Lip21d, Lip21h, Lip21i, Lip21j, Lip21k, Lip21l, Lip21m, Lip22c, Lip22d, Lip22a, Lip22b, Lip22e, Lip22l, Lip22f, Lip22g, Lip22h, Lip22i, Lip22k, Lip22j, Lip22m, Lip22n, Lip23a, Lip23b, Lip23f, Lip23c, Lip23d, Lip23e, Lip24a, Lip24b, Liu21, Liu23, Liu24, Lu20, Mal22a, Mal22b, Mal22c, Mal23a, Mal23b, Mal23c, Mal23d, Mal23e, MD24, Mal20, Mal21, Mar20, Mye21, Ndo24, Nel21]. Book [Oli20, Oli22, Rah23, Sa’23a, Sau21, Sau23, Sea20, SA23c, VL21a, VL21b, Wlu21, Wlu23, Zel22, ZGAR24]. Booth [Ano21b]. Bootstrap [Sea20]. Bordignon [Lip20f]. both [LCX23]. Boundary [QBN21]. Bowden [Gho20]. Bradley [Lip23b]. Bretz [APF24, Zel22]. Brooke [HL23]. Bruce [Ahm21d, Lip22h, Lip22n]. Burzykowski [Lip22f]. Business [Ano21a].

C [Liu23, Mal22a]. Caiado [VL21b]. Calibration [GBA^{+}21, GAM23, HG22, LT22]. Cambridge [Lip23b, Mal23d]. Cameron [Lip22h]. Can’t [Wu21]. Cantos [Lip20d]. Cantos-Gomez [Lip20d]. Cao [Lip22l]. Car [Liu24]. Carolyn [Mal22a]. Carroll [Lip22c]. Carrozzo [Lip20f]. Casino [Lip23g]. Causality [Kha23]. Ceja [Lip22e]. Celebrating [Jos21]. Censored [KTF21, KL20]. Censoring [LX23]. Century [Lip22h]. Cham [Ahm20a, Ahm20b, Lip20d, Lip20c, Lip20e, Lip20h, Lip20f, Lip20j, Lip22h, Lip24a, Lip24b, Mal20]. Chance [Mal22c, Lip23a, Lip23a]. Chang [Wlu21]. Change [DCW20, GYZ23, JYYJ22, WXX^{+}23, XWYL23, ZM23, ZHMY22, ZMXX23]. Change-Point [WXX^{+}23, XWYL23, ZM23, ZMXX23]. Changes [RK23]. Chapman [Ahm21d, Ano21e, AL22, APF24, GC23, Hus23, Kha23, Lip20a, Lip20b, Lip20g, Lip20i, Lip21c, Lip21g, Lip21f, Lip21e, Lip21d, Lip21h, Lip21i, Lip21j, Lip21k, Lip21m, Lip22c, Lip22d, Lip22l, Lip22f, Lip22i, Lip22k, Lip22j, Lip22m, Lip22n, Lip23f, Lip23c, Lip23d, Lip23g, Mal23b, Oli20, Oli22, Sau23, SA23c, VL21b, Zel22, Liu24]. Characteristics [LWHY21, ZY23]. Characterization [GMLB22]. Chart [CLM^{+}21, QLL20]. Charts [WT23]. Chatterjee [Mar20]. Chelsey [DM20]. Chen [Ahm21c, Ahm21e]. Chester [Lip20i]. Cheung [APF24, Zel22]. Choice [Lip22c, Lip22m, Mal22c]. Christian [Lip20h, Mal23a]. Christopher [Lip21i, Lip22c]. Claim [FYQ^{+}21, WZZ23]. Clark [Lip20g]. Class [KL20, RRH22]. Classification [LJ22, RRH22, VL21b]. Classifiers [WZ22]. Clinical [Ahm20a, Zel22]. Cloud [BA22]. Clustered [WZZ23]. Clustering [AX20, CJXD21, LK21, PHRD22, VL21b]. Codes [GSSI22]. Coefficient [WZZ23, WDWZ21]. Collaborators [Ano21f]. College [Ahm20c, Ano21d]. Collisions [LLK^{+}23]. Collocation [Lip20d]. Color [Lip23f, Lip21c, Lip22e, Lip22l, Lip22g, Lip22h, Lip24b, Liu23]. Columns [VGS20]. Comment [CSX^{+}21, FL21, GR21, LLBL21, Yua21, Zou21]. Company [DM20]. Comparison [PTD22]. Comparisons [Gho20]. Completion [SWL^{+}22]. Complex [STC^{+}22]. Component [CSX^{+}21, LL21, WV21, YSX21]. Component-Position [YSX21].

Components [MGG20]. Composite [DCH^{+}22]. Compound [LYW^{+}23]. Computational [LCL23b]. Computer [GSSI22, Gho21a, GBA^{+}21, HC21, KK21, LCX23, MMK^{+}22, SBD^{+}22, XJR23, Lip23b]. Concept [Has21, ZBA23]. Condition [LY23]. Conditional [CZ22]. Confidence [Dog21, WXJT21]. Conservation [VL21a]. Conservative [AGC^{+}21]. Constrained [DPJS21, PL21, WSSE21, WDWZ21, ZSPT21, YFL22]. Constraints [MZHW22]. Constructing [CPW24, NPM20]. Construction [JLM^{+}21]. Contemporary [Ahm21e]. Context [Hoe21]. Continuous [ZH21]. Control [CLM^{+}21, Che22, LXTP20, QLL20, QX22, WT23, ZM23, ZdC21, ZPS23]. Controlled [APF24, JB23, Zel22]. Convex [CJXD21]. Coordinates [PHRD22]. Copula [Lyu21]. Corain [Lip20f]. Cornel [Lip24a]. Corrected [PHRD22]. Correlated [LK21, QLL20, QX22, XQWL22, XQ24]. Correlation [RR21, WDZS22, YXW23, Lip23f]. Cost [YFL22]. Cost-constrained [YFL22]. Count [DCW20]. Count-Weighted [DCW20]. Counter [Lip23a]. Counter-Intuitive [Lip23a]. Course [Mal21, Sea20, WDWZ21]. Covariance [DHJH22, RK23, XSL^{+}23]. Covariate [DHJH22, LDY23, YQ21]. Covariate-Dependent [DHJH22]. Covariate-Regulated [LDY23]. Covariates [LWHY21, ZK22]. Coverage [LJK20, LZC21]. Craps [Lip23g]. CRC [Ahm21d, Ano21a, Ano21c, Ano21e, AL22, Asz23, DSB23, Els20, GC23, Hus23, Kha23, Lip20a, Lip20b, Lip20g, Lip20i, Lip21c, Lip21g, Lip21f, Lip21e, Lip21d, Lip21h, Lip21i, Lip21j, Lip21k, Lip21m, Lip22c, Lip22d, Lip22e, Lip22l, Lip22f, Lip22g, Lip22i, Lip22k, Lip22j, Lip22m, Lip22n, Lip23f, Lip23c, Lip23d, Lip23g, Mal22b, Mal23c, MD24, Mal21, Mye21, Oli20, Oli22, Sau23, Sea20, SA23c, VL21b, Wlu21, Wlu23, Zel22, Liu24]. CRC/Chapman [Lip22d, Zel22]. CRC/CRC [Lip22l, Lip23f]. Crellin [Lip20h]. Crime [GC23]. Criminologist [GC23]. Criteria [VE21]. Criterion [DSW23, HCL23]. Cross [DZ21]. Cross-Validation [DZ21]. Cruz [Lip21k]. Crystalline [MCD^{+}22]. Culture [Lip21l]. Curve [HPR^{+}21, PTD23, QXY20]. Curves [WZQ22]. Customer [Lip20f]. Cycle [Lip20j].

D [Ano21c, OKF22, SA23b]. D-Printed [OKF22]. Daniel [AL22]. Dark [Rah23]. Data [AS21, Ahm21b, Ahm21d, Ano21e, AL22, BA22, FYQ^{+}21, GYPS21, GLP22, GMLB22, GG22, HTLS21, Has21, JV22, KTF21, KL20, LJK20, LJ22, LCL23a, LBTM22, LXTP20, LX23, LPCX23, Lip20b, Lip20i, Lip21c, Lip21k, Lip22l, Lip22n, Lip23b, Lip23f, LP20, Liu24, LWHY21, LZYL22, MCD^{+}22, Mye21, PTD22, QLL20, QX22, QBN21, Rah23, RR21, Sau23, SWL^{+}22, TCXM23, WZQ22, WZZ23, XQWL22, XWYL23, YXW23, YA21, ZWX23, ZYLS21, ZM23, ZdC21, ZMXX23, ZGAR24, Ahm21d, DM20, Hus23, Lip20f, Lip22n, Nel21, Ahm21a, Els20]. Data-Driven [WZQ22]. Datasets [LP23]. Dattoli [Mal23e]. David [Cha22, Lip22c, Lip22a, Lip22b, Mal23a, Rah23]. Debates [Liu24].

Debunking [Lip21n]. Decision [Yan24, ZH21, Ahm20a]. Decision-Making [Ahm20a]. Decision-Oriented [Yan24]. Decomposition [HCL23, Yan24]. Deep [DSB23, MWG23, SGH23, YK23]. Defense [Lip21g]. Definitive [VGS20]. Deformable [DPJS21]. Deformation [QSK21]. Degradation [LWHY21, PC20, SYH21, WM21, ZY23]. Degree [PHRD22]. Degree-Corrected [PHRD22]. Delaigle [Lip23c]. Demand [XYWL21]. Demidenko [Lip21b]. Demonstration [WF21]. Density [CZ22, DPJS21, LJK20, LCL23a, LZYL22]. Density-Valued [LCL23a]. Dependence [WT23, ZK22]. Dependent [DHJH22, ZY23]. Depth [ZSD24]. Depth-based [ZSD24]. Depths [HTLS21]. Design [AGC^{+}21, ECNA20, Kha23, LCX23, Pou21, SBD^{+}22, YSX21, ZCG21, ZK22]. Designing [Ahm20b]. Designs [HS21, JLM^{+}21, KK21, MGG20, NPM20, NG20, SCA21, SS23, SAMJ23, VGS20, WGKJ24, WSSE21, XJR23]. Destination [XYWL21]. Destructive [SYH21]. Detecting [HTLS21, RK23]. Detection [DCW20, GMLB22, GYZ23, JYYJ22, LMN^{+}22, LCL23a, MF21, QBN21, TCXM23, WXX^{+}23, XWYL23, ZM23, ZHMY22, ZMXX23]. Determination [WZQ22, WF21]. Development [Wlu21]. Dever [Ahm20b]. Deviations [STC^{+}22]. Dexter [Ano21b]. Diagnosis [GLP22, XQWL22, XSL^{+}23]. Diagnostic [LXTP20, Ahm20a]. Diagnostics [ZBA23]. Diffusion [LY23]. Dimension [WY21]. Dimensional [CMJZ21, ECNA20, GLLL21, GLP22, GG22, Hus23, LZC21, LXTP20, MMK^{+}22, Mal23a, RR21, XQWL22, XSL^{+}23, YFL22, ZM23, XWYL23]. Dimensions [WSSR21]. Dimitris [Sea20]. Ding [Ahm21c, Ahm21e, Els20]. Ding-Geng [Ahm21c, Ahm21e]. Directional [Liu23]. Discontinuities [PGB^{+}21]. Discovery [Asz23, LXTP20]. Discrete [WXJT21]. Discrete-Event [WXJT21]. Discriminant [HSZK20]. Disease [YQ20]. Distance [Lip23f, WDZS22, ZCG21]. Distance-Distributed [ZCG21]. Distributed [SA23b, ZCG21]. Distribution [KTF21, LCL23a, VE21]. Distribution-to-Distribution [LCL23a]. Distributional [LCX23]. Dive [Lip20i]. Diversity [WDZS22]. Doganaksoy [Ano21a]. doi.org [DSB23]. Dolfini [Lip20h]. Domain [SA23b]. Donatello [Lip20g]. Don’t [Rah23]. Dorner [Mal23a]. Dr [Ndo24]. Drift [ZBA23]. Driven [WZQ22]. Driver [Liu24]. Dropping [VGS20]. Drug [Wlu21]. Dual [Tsa23]. Dual-Orthogonal [Tsa23]. D’Urso [VL21b]. Dylla [Lip21l]. Dynamic [GG22, LP20, MVK23, QXY20, XWYL23, ZYLS21]. Dynamical [LLK^{+}23].

ebk [Liu24]. eBook [Cha22, HL23, Lip21g, Lip21h, Mal22c, Wlu21, Mar20]. Ecology [Lip20c, VL21a]. Economists [Mal23b]. Ed [Ahm20a, Ahm20b, Ahm20c, AL22, Fir23, Lip20g, Lip21c, Lip22i, Lip23b, Lu20, Oli20, Sau21, SA23c]. ed. [Ahm21b, Lip23e]. ed.Edited [APF24]. Edited [Nel21, ZGAR24, Ahm21c, Ahm21d, Ahm21e, DSB23, Lip21c, Lip21g, Lip21f, Lip21e, Lip21d, Lip21k, Lip23c, Mal21]. Edition [Lip22c, Lip22j, Lip22n, Mal22a, Mal22b, Nel21, Ano21b]. Editor [Ano22c, Apl20, Jos23]. Editorial [Ano21f, Ano22c, Jos21, Pra22]. editors [Lip20a, Ano24]. eds [Zel22]. Education [Lip21a]. Effect [Kha23]. Effective [YQ20]. Effectiveness [YCTY24]. Effects [SS23, YPLS20, MF21]. Efficient [LH22, LZYL22, NPM20]. Efron [Lip23b]. Ejaz [Hus23]. Elastic [CHH^{+}23, HTLS21]. Elections [Lip21p]. Elements [Lyu21]. Elizabeth [VL21b]. Ellabban [Ndo24]. Emulating [HC21]. Emulation [MMK^{+}22, MWG23]. Emulator [LT22]. Enabling [Ndo24]. Encarnación [Lip20a]. Energy [DSW23, Lip23f, Els20]. Engineering [Ano21b, ECNA20]. Enrique [Lip22e]. Ensuring [MGG20]. Enumeration [NG20]. Envelopes [AS21]. Environment [LP20]. ePDF [Nel21]. Epidemiology [Mal22b]. ePub [Nel21]. Equation [Lip21o]. Error [LJK20, Lip21h, Lip23c, MCD^{+}22, MGG20]. Errors [DZ21]. Escobar [Lu20, Nel21]. ESD [Che22]. Essential [Has21]. Estimating [LY23]. Estimation [AGC^{+}21, CMAKA23, DPJS21, LJK20, LZYL22, MGG20, QSK21, WY21, XKSK21, ZSD24]. Estimator [SGG^{+}22]. Ethics [MD24]. Eugene [Lip21b]. EUR [HL23, Mal22c, Sa’23a]. Evaluating [QXY20]. Evaluation [VGS20]. Event [KL20, WV23, WXJT21, YA21]. Events [WWYC24]. Ever [Mal22a]. Evidence [Lip20c, Lip23b]. Exact [WZ22]. Examine [Lip22f]. Examples [Lip22g, Mal21]. Exciting [WXX^{+}23]. Excursion [AGC^{+}21]. Experimental [KK21, MZHW22]. Experimentation [MZHW22]. Experiments [AGC^{+}21, ECNA20, HS21, JB23, LCX23, LC23, MMK^{+}22, RD24, SCA21, SGG^{+}22, SKD20, SBD^{+}22, Tsa23, WDWZ21, XJR23, YSX21]. Experts [MVK23]. Explain [Lip22f]. Explanatory [Lip22f]. Explore [Lip22f]. Explores [Lip21l]. Exponentially [Dix22]. Extendable [LH22]. Extremely [TCXM23].

Fabrication [WDWZ21]. Facial [ZSPT21]. Factor [XJR23]. Factorial [HS21, Tsa23]. Factors [ZTCA21]. Failure [WWYC24]. Failures [LBTM22]. Fans [Fir23]. Faraway [Lip21j]. Faris [Liu24]. Fast [RR21, WZ22, ZSD24]. Fault [LDY23, XQWL22]. Feature [FL21, GG22, HC21, TCXM23]. Features [KK21, WDZS22]. Federated [CA24, SA23b]. Feedback [ZPS23]. Fei [DSB23]. Fernandes [Liu21]. Feryaal [Hus23]. Fidelity [SBD^{+}22]. Field [LBTM22, Ahm21e]. Filling [WGKJ24]. Filter [PL21]. Filtering [GMLB22]. Finch [Lip22d]. Finite [ZH21]. First [Sea20]. Fisher [Yan24]. Fitting [LH22]. FL [Ahm21d, Ano21a, Ano21e, AL22, Asz23, DSB23, Els20, GC23, Lip20a, Lip20b, Lip20g, Lip20i, Lip21c, Lip21g, Lip21f, Lip21e, Lip21d, Lip21h, Lip21i, Lip21j, Lip21k, Lip21m, Lip22c, Lip22d, Lip22e, Lip22l, Lip22f, Lip22g, Lip22i, Lip22k, Lip22j, Lip22m, Lip22n, Lip23f, Lip23c, Lip23d, Lip23g, MD24, Mal21, Mye21, Oli20, Oli22, Sau23, Sea20, SA23c, VL21b, Wlu21, Wlu23, Zel22]. Florence [Liu23]. Fog [CKJD23]. Foldover [NPM20]. Football [Ahm21b]. Forecasting [Dix22]. Foreword [ZGAR24]. Formulation [Dog21]. Foster [Lip21c]. Foundations [Sau23]. Four [Lip23g]. Fourier [HC21]. Fragnelli [Lip20a]. Framework [MZHW22, XQ24, YSM23]. Francis [Ahm21d, Ano21a, Ano21c, Ano21e, Asz23, Els20, GC23, Lip20a, Lip20b, Lip20g, Lip20i, Lip21c, Lip21g, Lip21f, Lip21e, Lip21d, Lip21h, Lip21i, Lip21j, Lip21k, Lip21m, Lip22c, Lip22g, Lip22m, Lip22n, Lip23f, Lip23c, Lip23d, Lip23g, Mal22b, Mal21, Mye21, Nel21, Oli20, Sea20, Wlu21, Lip22d, Lip22e, Lip22l, Lip22f, Lip22i, Lip22k, Lip22j]. Francisco [Lip21k]. Frand [Zel22]. Frank [APF24]. Frauke [Ahm20b, Lip21c]. Frede [Ndo24]. Frederick [Lip21l]. Fridley [HL23]. Frontiers [Ahm21c]. Function [CMJZ21, HLPK21, PTD22, SBWK22, Lip21h]. Function-on-Function [CMJZ21, HLPK21, SBWK22]. Functional [AS21, CLM^{+}21, CHH^{+}23, DHJH22, HTLS21, LCL23a, LT22, MMK^{+}22, ZYLS21]. Fused [CHH^{+}23]. Fuselage [DCH^{+}22]. Future [LBTM22].

G [Ano21a, MD24, Nel21]. Gabor [Lip23f]. Gamblers [Fir23]. Games [Lip23g, Lip23e]. Gamma [KTF21]. Gap [DCH^{+}22]. Garcia [Lip22e]. Gaussian [CKL21, CA24, LP23, LT22, LPCX23, LJ21, MWG23, PGB^{+}21, Pou21, PTD22, SGH23, ZCG21, ZTCA21]. Gauvreau [Wlu23]. GBP [DSB23]. General [LWHY21, XQ24, YSM23]. Generalized [CJXD21, GBA^{+}21, GG21, KTF21, LL21]. Generative [YK23]. Geneva [Ahm20d]. Geng [Ahm21c, Ahm21e]. Genomics [HL23]. Geometrical [STC^{+}22, ZdC21]. Geometry [GSSI22, Mal23a]. Geophysical [GAM23]. Gerald [Ano21a, Lu20]. Geretschläger [Mal23a]. Ghani [Lip21c]. Gilbert [Ahm20d]. Giuseppe [Ahm20a, Ano21e, Mal23e]. Globe [Ano21b]. Gomez [Lip20d]. Grace [Lip23c]. Gradient [HC21]. Graduate [SA23c]. Graduate-Level [SA23c]. Graphical [WZL22]. Great [Liu24]. Grid [Ndo24]. Group [Ahm21d, Ano21a, Ano21c, Ano21e, Asz23, Els20, GC23, JLM^{+}21, Lip20a, Lip20b, Lip20g, Lip20i, Lip21c, Lip21g, Lip21f, Lip21e, Lip21d, Lip21h, Lip21i, Lip21j, Lip21k, Lip21m, Lip22c, Lip22d, Lip22e, Lip22l, Lip22f, Lip22g, Lip22i, Lip22k, Lip22j, Lip22m, Lip22n, Lip23f, Lip23c, Lip23d, Lip23g, Mal22b, Oli20, Sea20, WSSE21, MF21]. group-based [MF21]. Group-Orthogonal [JLM^{+}21]. Guanquan [Mye21]. Guenter [Mal23a]. Guide [GC23, Lu20, Lip21k, Mal22a, Mal22b, MD24]. Gustafson [Lip23c].

H [Asz23, Lip21l, Mal20]. Hadamard [NPM20]. Hahn [Ano21a, Lu20]. Haining [Mye21]. Haitham [Ndo24]. Hall [Ahm21d, Ano21e, AL22, GC23, Hus23, Kha23, Lip20a, Lip20b, Lip20g, Lip20i, Lip21c, Lip21g, Lip21f, Lip21e, Lip21d, Lip21h, Lip21i, Lip21j, Lip21k, Lip21m, Lip22c, Lip22l, Lip22f, Lip22i, Lip22k, Lip22j, Lip22m, Lip22n, Lip23f, Lip23c, Lip23d, Lip23g, Mal23b, Oli20, Oli22, Sau23, SA23c, VL21b, APF24, Lip22d, Liu24, Zel22]. Hall/CRC [Ano21e, AL22, GC23, Hus23, Kha23, Lip20a, Lip20b, Lip20g, Lip20i, Lip21c, Lip21g, Lip21f, Lip21e, Lip21d, Lip21h, Lip21i, Lip21j, Lip21k, Lip21m, Lip22f, Lip22i, Lip22j, Lip22n, Lip23c, Lip23d, Oli20, Sau23, SA23c, VL21b, Liu24].

Hall/Taylor [Lip23f]. Hampson [APF24, Zel22]. Hand [Rah23]. Handbook [Ahm21a, Che22, APF24, Lip20a, Lip21g, Lip21f, Lip21e, Lip21d, Lip22g, Lip23c, Zel22]. Hardback [Ano21e, AL22, DSB23, Lip21b, Lip21c, Lip21j, Lip21k, Lip21m, Mal21, Nel21, VL21b, Lip21g, Lip21h, Mal22c, Wlu21]. hardcover [Liu24]. Hare [Lip22c]. Harezlak [Cha22]. Hastie [Lip23b]. Hayward [SA23c]. HB [Mal22b, Mal23a, Mal23b, Mal23e]. hbk [Che22, Lip22c, Lip22d, Lip22l, Lip22f, Lip22g, Lip22h, Lip22j, Lip23f, Lip23c, Lip24a, Liu23, Liu24, SA23c]. hcb [Mal23c, Sau23]. Healthcare [Wlu21]. Heather [Wlu23]. Hei [DSB23]. Henson [SA23c]. Heredity [SKD20]. Heterogeneity [PGB^{+}21]. Heterogeneous [GYPS21, LBTM22, LP20, WZL22]. heteroscedasticity [MF21]. Hierarchical [JYYJ22, LH22, QBN21]. Hierarchy [Lip21m]. High [GLLL21, GLP22, GG22, Hus23, LXTP20, LZYL22, MMK^{+}22, RR21, XQWL22, XSL^{+}23, XWYL23, YFL22, ZM23]. High-Dimensional [GLLL21, GLP22, GG22, LXTP20, MMK^{+}22, RR21, XQWL22, XSL^{+}23, YFL22, ZM23, XWYL23]. High-Throughput [LZYL22]. Hilbert [LYW^{+}23]. Hill [DM20]. Historical [Hoe21]. History [Lip21l]. Hoboken [Lip21b, Lu20, Ndo24]. Hockey [ZGAR24]. Holmes [Lip22d]. Honeycomb [HG22]. Horn [Lip20h]. Horton [AL22]. Howard [Lip21g, Lip22c]. https [DSB23]. Hu [DSB23]. Hubele [Liu24]. Human [Mal22c]. Humana [HL23]. Humanities [Lip20j]. Hygiene [ABRA20]. Hypothesis [TCXM23].

Iacus [Ano21e]. Ian [Lip21c]. Ideas [Lip20j]. IEEE [Che22, Ndo24]. II [Lip21g, Lip22c]. ill [Lip21c, Lip22d]. Illustrations [DSB23, Lip23f, Lip21i, Lip21j, Lip21k, Lip22e, Lip22l, Lip22g, Lip22h, Lip22k, Lip23c, Lip23g, Lip24b, Liu23, Liu24]. Image [SWL^{+}22, ZPS23]. Image-Based [ZPS23]. Images [MCD^{+}22, TCXM23]. Imaging [MCD^{+}22]. Imperfect [GAM23]. Important [WSSR21]. Impossible [Lip24a]. Improved [CSX^{+}21]. Imputation [MWG23]. Incidence [Mal23d]. Incorporated [DM20]. Incorrectly [LJ22]. Individual [LJ22]. Industrial [ABRA20, Dix22]. Industry [Pra22, Lip24b]. Inference [Asz23, Lip20i, Lip23b, RD24]. Information [GSSI22, Gho20, HC21, MD24, YQ21, Yan24]. Initial [LY23]. Initialization [WXJT21]. Inkjet [LZC21]. Innovative [Ahm20d, Liu23]. Input [LT22]. Inputs [LCX23, ZH21]. Inspection [YPLS20]. Inspiring [Zou21]. Insurance [FYQ^{+}21]. Integrals [Lip24a, Lip21h]. Integration [Lip22l]. Intelligence [Sa’23a, Wlu21]. Intensive [Gho21a]. Interferograms [GAM23]. International [Ahm20a, Ano21b, Lip21l, Liu21]. International/Red [Ano21b]. Interpolation [Xio21]. Intervals [Dog21, Lu20]. Intrinsic [ZdC21]. Introduction [Kha23, Lip22i, Lip23a, Oli22, Lip20b, Lip23d]. Intuition [Lip23a]. Intuitive [Lip23a]. Inverse [LY23, WY21]. Investing [Lip24b]. Ioan [Lip24a]. Irizarry [Lip20b]. ISBN [Ahm20a, Ahm20b, Ahm20c, Ahm20d, Ahm21c, Ahm21d, Ahm21e, Ano21a, Ano21b, Ano21c, Ano21d, Ano21e, AL22, APF24, Asz23, Cha22, Che22, DM20, DSB23, Els20, Fir23, GC23, Gho20, HL23, Hus23, Kha23, Lip20a, Lip20b, Lip20d, Lip20c, Lip20e, Lip20h, Lip20f, Lip20g, Lip20i, Lip20j, Lip21b, Lip21c, Lip21g, Lip21f, Lip21e, Lip21d, Lip21h, Lip21i, Lip21j, Lip21k, Lip21l, Lip21m, Lip22c, Lip22d, Lip22a, Lip22b, Lip22e, Lip22l, Lip22f, Lip22g, Lip22h, Lip22i, Lip22k, Lip22j, Lip22m, Lip22n, Lip23a, Lip23b, Lip23f, Lip23c, Lip23d, Lip23g, Lip24a, Lip24b, Liu21, Liu23, Lu20, Mal22a, Mal22b, Mal22c, Mal23a, Mal23b, Mal23c, Mal23d, Mal23e, MD24, Mal20, Mal21, Mar20, Mye21, Ndo24, Nel21, Oli20, Oli22, Rah23]. ISBN [Sa’23a, Sea20, SA23c, VL21b, Wlu21, Wlu23, Zel22, ZGAR24, Liu24]. Ishii [Mal22c]. Ismay [Lip20i]. Isolation [LDY23, ZHMY22]. Issue [Pra22]. Item [Lip21f, Lip21e, Lip21d]. Iterative [FL21]. ix [Ahm20c, Ano21d, Lip20d].

J [Ano21b, AL22, Lip20h, Lip21f, Lip21e, Lip21d, Lip23f, Lu20, Mal23b, Oli20, Rah23]. J. [Lip21j]. Jacob [GC23]. James [Lip21g, SA23c]. Jarmin [Lip21c]. Jaroslaw [Cha22]. Jeremy [Che22]. Jessop [Lip20c]. Jill [Ahm20b]. Joaquín [Lip20a]. John [Lip21b, ZGAR24]. Joint [LK21, LRWD23, XSL^{+}23, YA21]. Jorge [VL21b]. Joseph [Ano22c]. Jossey [Mal22a]. Jossey-Bass [Mal22a]. Journeys [Lip21l]. JST [LRWD23]. JST-RR [LRWD23]. Judgment [Lip22c]. Julia [Lip21c]. Julian [Lip21j]. Jumps [WZQ22].

K.A [Ano21b]. Kalajdzievski [Lip22i]. Kalpana [Ano21c]. Kao [Asz23]. Kao-Tai [Asz23]. Kaplan [AL22, GC23]. Kateri [Sau23]. Kay [Mal23c]. Keith [Lip22c, Lip22g]. Kernel [LYW^{+}23]. Kevin [ZGAR24]. Key [Ano21a]. Kim [APF24, Lip20i, Lip22l, Zel22]. Kim-Ahn [Lip22l]. Kimberlee [Wlu23]. Kindle [Nel21]. Klaas [Lip22k]. Know [Rah23]. Knowledge [Asz23, Lip21o]. Konrad [Lip21m]. Konstantinos [Fir23]. Korosteleva [Oli22]. Korotkov [Lip21h]. Kreuter [Ahm20b, Lip21c]. Kriging [CMJZ21]. Kuen [APF24, Zel22]. Kulakowski [Lip21m]. Kumiko [Mal22c]. Kyung [Zel22]. KyungMann [APF24].

L [Fir23, Lip20f, Lip22k, Lip23f, Mal20, SA23c, HCL23]. Lê [Lip22l]. Label [LJ22, YK23]. Label-Noise [YK23]. Labeled [LJ22]. Lane [Lip21c]. Language [Mal22c]. Large [LLBL21, LMN^{+}22, LPCX23, LP20, MMK^{+}22, WZ22]. Large-Margin [WZ22]. Large-Scale [LMN^{+}22, LPCX23, MMK^{+}22]. Lasso [DZ21]. Latent [LMN^{+}22, ZTCA21, MF21]. Law [Lip20c]. Lawrence [Ahm20c, Ano21d]. Laws [CKL21]. Layer [LZC21]. Learning [Ahm21d, Asz23, CKL21, DSB23, DCH^{+}22, Hus23, Lip22e, Lip22n, Liu21, LDY23, LLK^{+}23, QX22, SGH23, SA23b, XWYL23, YK23, ZYLS21, ZBA23, Liu21, Lip21i]. Least [WV21]. Leave [WZ22]. Leave-One-Out [WZ22]. Leemis [Ahm20c, Ano21d]. Leighton [Lip22m]. Let [Lip20c]. Level [NPM20, SS23, SA23c, Tsa23]. Lexical [Lip20d]. Li [Mye21]. Licciardi [Mal23e]. Lies [Lip23e]. Life [Lip20j]. Lifetime [KTF21, LX23]. Lifetime-Censored [KTF21]. Likelihood [CJXD21, Dog21]. Linden [Lip21f, Lip21e, Lip21d]. Linear [Ano21c, CJXD21, HSZK20, MF21, Ano21c, Lip21j]. Lingjie [Lip24b]. Lisa [APF24, Zel22]. LLC [Asz23]. Local [KK21, WBG22]. Locally [WSSR21, ZK22]. Location [MCD^{+}22, ZSD24]. Logic [Lip23a, Lip23e]. Logistic [LJ22]. London [Ano21b]. Long [SGG^{+}22]. Long-Term [SGG^{+}22]. Low [HCL23]. Low-Rank [HCL23]. Luck [Lip23e]. Luis [Lu20, Nel21].

M [Ahm20c, Ano21d, Ano21e, Che22, Lip21g, Ndo24, SA23c]. Ma [Lip24b]. Machine [Ahm21d, DSB23, Hus23, Lip22e, Lip22n, Asz23, Liu21]. Machine-Learning [Ahm21d, Lip22n]. Macmillan [Ano21b]. Maharaj [VL21b]. Main [SS23]. Maintenance [YCTY24]. Mair [Lip20e]. Making [Ahm20a]. Management [Ahm21a]. Managers [Fir23]. Manhattan [Lip22h]. Manifold [ZdC21]. Mann [Zel22]. Manufacturing [CKJD23, STC^{+}22, SWL^{+}22]. Maps [RRH22]. Marcello [Mal23e, Wlu23]. Maresch [Mal23a]. Margin [WZ22]. Maria [Lip23f, Sau23]. Marion [Lip20h]. Mark [Lip23g, Wlu21]. Markov [CMAKA23]. Mary [Ahm20c, Ano21d]. Mass [WWYC24]. Massive [LP23]. Materials [LZC21, MCD^{+}22]. Math [Lip23d, Mal22a, ZGAR24, Lip22i]. Mathematical [Mal22c, Mal23e, Ano21d, Lip22j, Mal23b, SA23c]. Mathematics [Ahm21b, Ano21b, Fir23, Lip22a, Lip23e, Mal23e, SA23c, Ano21b, Lip22b, Lip22i, Mal23c, Lip23g]. Mathletics [Fir23]. Matrices [LLBL21, NPM20]. Matrix [HSZK20, RK23, XSL^{+}23]. Matt [Cha22]. Matters [Rah23]. Mattias [Ahm20d]. Mattie [Wlu23]. Maximin [VE21]. Maximizing [DM20, SBD^{+}22]. Maximum [DCH^{+}22, XJR23]. May [Lip20g]. Mayer [Mal20]. McElreath [Mal21]. McElroy [Sea20]. McNulty [Lip22g]. Mean [DHJH22, WY21, XSL^{+}23]. Measurement [DZ21, Gho21b, GAM23, LJK20, Lip23c, LP20, MCD^{+}22, Lip22k]. Measures [WM21]. Media [Ano21e]. Medicine [Lip20c, Wlu21]. Meeker [Ano21a, Lu20, Nel21]. Mello [Liu21]. Meta [Ahm20a]. Meta-Analysis [Ahm20a]. Method [DSW23, HC21, JV22, YQ20, ZSD24, ZMXX23]. Methodologies [Asz23]. Methods [Ahm20d, Ahm21b, APF24, Gho21a, Lip21c, Lip22d, Lip22l, Mal23d, Mar20, Nel21, QXY20, SA23c, WM21, Yua21, Zel22]. Metrics [CKJD23]. Metrology [OKF22]. Michael [Ahm21a, Lip23d, Mal23b]. Microeconometrics [Lip21i]. Military [Lip21g]. Minimally [NG20]. Minimum [DSW23, WXJT21]. Mining [Ahm21d, Lip22n]. Miroslav [Ndo24]. Missed [LXTP20]. Misspecification [KTF21]. Mitchell [Ahm21a]. Mixed [CKJD23, NPM20, YPLS20]. Mixed-Level [NPM20]. Mixing [GR21]. mixOmics [Lip22l]. Mixture [LJK20, LK21, MVK23, SKD20]. Mixture-of-Mixtures [SKD20]. Mixtures [SKD20]. Moacir [Liu21]. Mobile [LJK20]. Modality [DPJS21]. Modality-Constrained [DPJS21]. Model [AX20, DCH^{+}22, GBA^{+}21, JYYJ22, KTF21, KK21, LL21, LJK20, LK21, LP23, LYW^{+}23, LRWD23, Lip22f, LDY23, MMK^{+}22, QBN21, SKD20, SBWK22, WDWZ21, WZL22, XSL^{+}23, YSX21, YK23, YPLS20, ZY23, ZH21, MF21]. Model-Based [AX20, ZH21]. Modeling [ABRA20, Ahm21e, Asz23, DCW20, FYQ^{+}21, GYPS21, Gho21c, LZC21, LRWD23, Lip22g, Lip22j, Lip22n, Lyu21, PGB^{+}21, Pou21, PTD23, RD24, STC^{+}22, SYH21, SA23c, VL21a, WBG22, XYWL21, XKSK21, YCTY24, YSM23, ZYLS21, ZTCA21]. Modelling [Ahm21d, Mye21]. Models [CMAKA23, CJXD21, GLLL21, GAM23, HC21, KK21, LMN^{+}22, LH22, Lip21j, Lip22c, Lip22k, Lip22j, LY23, LWHY21, MVK23, SS23, WZZ23, YA21, ZBA23, MF21, Lip21f, Lip22f, Lip23c]. Modern [Lip20i, Mal23e, AL22, Lip20e]. Modification [VE21]. Moises [Lip20d]. Monitoring [GLP22, LX23, LDY23, MMO24, QLL20, QXY20, STC^{+}22, WT23, WZL22, XQ24, YQ21, YXW23, YQ20, ZWX23, ZM23, ZBA23]. Mukherjee [Mar20]. Multi [CKJD23, CA24, HG22, LJK20, SCA21, SWL^{+}22, SBD^{+}22, YXW23]. Multi-Accuracy [LJK20]. Multi-Coverage [LJK20]. Multi-Fidelity [SBD^{+}22]. Multi-Output [CA24, HG22]. Multi-Platform [SCA21]. Multi-Profile [YXW23]. Multi-Responses [CKJD23]. Multi-Sources [SWL^{+}22]. Multicriteria [NG20]. Multifidelity [KK21, SBWK22]. Multilayer [DCW20]. Multiple [GYPS21, GAM23, LWHY21, WV23]. Multiscale [QBN21]. Multisource [FYQ^{+}21]. Multivariate [ECNA20, JYYJ22, Lip20g, SYH21, WXJT21, WZL22, XQ24, ZY23, ZYLS21, Lip22l]. Murugan [DM20]. Mutually [WXX^{+}23]. Myths [Lip21n].

N [Ahm21a, Lip21h, Sea20]. Nanomanufacturing [YPLS20]. Nanoparticle [TCXM23]. Natalie [Lip21g]. Nature [Ahm21c, Ahm21e, Lip21l, Lip22h, Lip24a, Lip24b, Liu21, Mar20]. Necip [Ano21a]. Need [Mal22a]. Nestler [Fir23]. Net [CHH^{+}23]. Network [YSM23, ZK22]. Networks [AX20, Dix22, DCW20, LMN^{+}22, MMO24]. Neural [Dix22, MMO24]. Nicholas [AL22]. Nightingale [Liu23]. Nikolai [Lip21h]. NJ [Lip21b, Lu20, Ndo24, Rah23]. Noise [LJ22, YK23]. Noisy [PTD22, TCXM23]. Nolan [DM20]. Non [Ano21c, LP23]. Non-Gaussian [LP23]. Non-Linear [Ano21c]. Nonconvex [YFL22]. Nonlinear [CHH^{+}23, LLK^{+}23]. Nonnested [KK21]. Nonparametric [AX20, BA22, GLLL21, LJK20, LK21, Lip20f, Mal20, WT23, XKSK21, ZWX23]. Nonstationary [LP23, QSK21]. Norma [Liu24]. Note [DZ21]. Novel [Dog21]. Novelty [SGG^{+}22]. Number [WZQ22, Mal23c]. Numbers [GC23]. Numeral [LCX23]. NY [DM20, HL23, ZGAR24].

O [Oli22]. Observable [GYZ23]. Observed [CPW24, ZHMY22]. Observing [MMK^{+}22]. Olympiad [Lip23d]. Omar [Ndo24]. On-site [HG22]. One [RD24, WZ22, XJR23]. One-Factor-At-A-Time [XJR23]. One-Shot [RD24]. Online [GLP22, GYZ23, HS21, JB23, MZHW22, MVK23, SCA21, SGG^{+}22, XWYL23, YQ20, ZWX23, ZMXX23]. Operations [Lip21g].

Optimal [JB23, JV22, MGG20, SAMJ23, ZK22]. Optimization [HSSP22, Mal23b, PL21, Pou21, WSSR21, YFL22, ZH21]. Order [LC23, Tsa23, YSX21, ZSPT21]. Order-Constrained [ZSPT21]. Order-of-Addition [LC23, Tsa23, YSX21]. Ordinal [WT23]. Oriented [Yan24]. Origin [XYWL21]. Origin-Destination [XYWL21]. Orthogonal [JLM^{+}21, NG20, Tsa23, WSSE21]. Other [Ahm21b, Lip20c]. Our [Mal23a]. Outlier [LCL23a]. Output [CPW24, CA24, HG22, MMK^{+}22, WGKJ24]. Overfitting [PTD23].

P [Cha22, Lip20f, Lip21g, Lip21i, Mar20]. Package [Lip22l]. Pagano [Wlu23]. pages [Liu24]. Panik [Mal23b]. Paperback [Lip21f, Lip21e, Lip21d, Lip21l, Ano21b, APF24]. Parameter [Yan24]. Parameters [LL21]. Parametric [LZYL22, Mal20, Lip20f]. Pareto [HSSP22]. Part [BA22, Liu21]. Part-to-Part [BA22]. Partially [CPW24, GYZ23, ZHMY22]. Partitioned [LP23]. Partitioning [GG22]. Pascual [Lip20d, Nel21]. Path [LWHY21, Mal23c]. Patrick [Lip20e]. Patterns [WT23]. Paul [Lip23c]. PBK [Mal22a, Kha23, Lip22a, Lip22b, Lip22i, Lip22k, Lip22m, Lip22n, Lip23a, Lip23b, Lip23d, Lip23g, Lip24b, Liu24]. PCA [DHJH22]. Pelechrinis [Fir23]. Penalized [CMAKA23]. Penalty [GR21]. People [Lip22g]. PERCEPT [ZMXX23]. Perform [HSSP22, Wu21]. Performance [ZY23]. Periodic [LPCX23]. Permutation [Gho21b]. Personalities [Lip21l]. Personalized [SA23b]. Perspective [DSB23]. Phi [VE21]. Philip [Mal20]. Philosophy [Lip21o]. Physical [ABRA20, CKL21]. Physicist [Lip21l]. PICAR [LH22]. Pierpaolo [VL21b]. Planning [WM21]. Platform [SCA21]. Plot [MGG20]. Point [BA22, JYYJ22, WXX^{+}23, XWYL23, ZM23, ZMXX23]. Points [CZ22, VE21]. Poisson [LYW^{+}23, WZZ23]. Political [Lip21k]. Politis [Sea20]. Polling [Lip21p]. Pollution [AX20]. Polynomial [Mal23d]. Pontintor [Liu21]. Poole [Lip22c]. Porro [Ano21e]. Posamentier [Mal23a]. Position [YSX21]. Positioning [LJK20]. Post [XQWL22, Hus23]. Post-Shrinkage [Hus23]. Post-Signal [XQWL22]. Power [PTD23]. pp [Ahm20a, Ahm20b, Ahm20c, Ahm20d, Ahm21c, Ahm21d, Ahm21e, Ano21a, Ano21b, Ano21c, Ano21d, Ano21e, AL22, APF24, Asz23, Cha22, Che22, DM20, DSB23, Els20, Fir23, GC23, Gho20, HL23, Hus23, Kha23, Lip20a, Lip20b, Lip20d, Lip20c, Lip20e, Lip20h, Lip20f, Lip20g, Lip20i, Lip20j, Lip21b, Lip21c, Lip21g, Lip21f, Lip21e, Lip21d, Lip21h, Lip21i, Lip21j, Lip21k, Lip21l, Lip21m, Lip22c, Lip22d, Lip22a, Lip22b, Lip22e, Lip22l, Lip22f, Lip22g, Lip22h, Lip22i, Lip22k, Lip22j, Lip22m, Lip22n, Lip23a, Lip23b, Lip23f, Lip23c, Lip23d, Lip23g, Lip24a, Lip24b, Liu21, Liu23, Lu20, Mal22a, Mal22b, Mal22c, Mal23a, Mal23b, Mal23c, Mal23d, Mal23e, MD24, Mal20, Mal21, Mar20, Mye21, Ndo24, Nel21, Oli20, Oli22, Rah23]. pp [Sa’23a, Sau23, Sea20, SA23c, VL21b, Wlu21, Wlu23, Zel22, ZGAR24]. ppb [ZGAR24]. Practical [Ahm21a, Lip21k, Liu21, Mal22b, MD24, Ahm20b, DM20, Lip20g]. Practice [Lip21c]. Practitioners [Lu20]. Precise [MGG20]. Precision [Wlu21]. Prediction [CKJD23, CJXD21, LBTM22, LRWD23, Lip20b, Lip21p, MVK23, WWYC24, XYWL21, YA21, Asz23]. Predictive [Ahm21d, Lip22f, Lip22n]. Prehistoric [Lip20h]. Presence [ZK22]. Press [Ano21b, Ano21c, AL22, Asz23, Che22, DSB23, Els20, Fir23, GC23, Hus23, Lip21h, Lip22c, Lip22e, Lip22l, Lip22g, Lip22i, Lip22k, Lip22j, Lip22m, Lip23b, Lip23f, Lip23c, Lip23d, Lip23g, Mal22b, Mal23c, Mal23d, MD24, Mye21, Ndo24, Oli22, Rah23, Sea20, SA23c, VL21b, Wlu23]. Press/Chapman [Lip22m, Lip23f, Lip23g]. Preventive [YCTY24]. Prewarping [WBG22]. Primacy [SGG^{+}22]. Princeton [Fir23, Rah23]. Principal [CSX^{+}21, LL21, WV21]. Principles [Wlu23]. Printed [OKF22]. Printing [CMJZ21, LZC21, SA23b]. Prior [CHH^{+}23]. Priors [HPR^{+}21]. Prizes [Ano20a, Ano22a]. Probability [Gho21c, Ahm20c, Lip22m]. Problem [PTD23]. Problems [DSW23, Lip21a, Lip23d]. Procedure [LXTP20]. Process [CKL21, LT22, LPCX23, LDY23, MMO24, MWG23, PGB^{+}21, Pou21, PTD22, QLL20, QXY20, QX22, SGH23, WZZ23, XSL^{+}23, YQ21, ZCG21, ZTCA21, ZdC21, Lip21m]. Process-Aided [PTD22]. Processes [ABRA20, CA24, GYPS21, LZC21, LJ21, LY23, Oli22, PC20, QSK21, WXX^{+}23, XKSK21, XQ24, YSM23]. Product [Ano21a]. Profile [YXW23]. Profiles [WZL22]. Program [Che22]. Programming [Oli20]. Project [Lip22h]. Projection [LL21]. Projections [VGS20]. Promote [WDZS22]. Properties [JLM^{+}21]. Property [Dog21]. Przemyslaw [Lip22f]. Psychological [Lip22k]. Psychometrics [Lip20e]. Publisher [Lip21l]. Publishing [Ahm20a, DM20, Liu21, ZGAR24]. Purchase [CJXD21]. Pure [MGG20]. Pure-Error [MGG20]. Python [Lip21j, Lip22g, Sa’23a, Sau23].

Q [Ano21a, Lu20, Nel21]. Qualitative [Lip20h, ZTCA21]. Quantification [GLLL21, GSSI22]. Quantify [HSSP22]. Quantile [AS21]. Quantitative [Lip20h, ZTCA21, Lip24b]. Quick [LC23]. Quick-Sort [LC23].

R [Asz23, GC23, Lip20f, Lip20i, Lip22l, Lip22g, Mal21, Oli22, Sau23, Lyu21, AL22, Cha22, Lip20b, Lip20e, Lip21b, Lip21i, Lip22e, Lip21k, Sau21, VL21a]. Rachel [Lip20h]. Radiation [GBA^{+}21]. Radin [Lip23d]. Radiosonde [AS21]. Rafael [Lip20b]. Random [HC21, RK23]. Randomized [APF24, Zel22]. Range [QLL20]. Rank [HCL23, ZWX23]. Rank-Based [ZWX23]. Ranking [CSX^{+}21, WDZS22]. Rate [LXTP20, SBD^{+}22]. Ratings [LRWD23]. Ratio [Dog21]. Ratner [Ahm21d, Lip22n]. Raton [Ahm21d, Ano21a, Ano21e, AL22, Asz23, DSB23, Els20, GC23, Lip20a, Lip20b, Lip20g, Lip20i, Lip21c, Lip21g, Lip21f, Lip21e, Lip21d, Lip21h, Lip21i, Lip21j, Lip21k, Lip21m, Lip22c, Lip22d, Lip22e, Lip22l, Lip22f, Lip22g, Lip22i, Lip22k, Lip22j, Lip22m, Lip22n, Lip23f, Lip23c, Lip23d, Lip23g, MD24, Mal21, Mye21, Oli20, Oli22, Sau23, Sea20, SA23c, VL21b, Wlu21, Wlu23, Zel22]. Rayid [Lip21c]. readers [Nel21]. Real [XYWL21, ZM23]. Real-Time [XYWL21, ZM23]. Reason [Lip22m]. Recognition [ZSPT21]. Reconciling [Ano21e]. Reconstruction [Xio21]. Recreational [Lip22a, Lip22b]. Recurrent [Dix22, KL20]. Red [Ano21b]. Reduction [DCH^{+}22, JB23, SBD^{+}22, WY21]. Reed [Lip22h]. Refaat [Ndo24]. Regime [CMAKA23]. Regime-Switching [CMAKA23]. Region [WXJT21]. Registration [HPR^{+}21]. Regression [CLM^{+}21, Cha22, CZ22, CHH^{+}23, CKJD23, CSX^{+}21, CLSZ21, DSW23, FL21, GYPS21, GG21, GG22, Hoe21, HLPK21, Jos21, LLBL21, LJ22, LCL23a, Lip22g, LCL23b, WZQ22, WZZ23, Wu21, Xio21, YFL22, Yua21, ZSPT21, Zou21, MF21]. Regressive [CMAKA23]. Regularization [GR21, Has21, LLBL21, Lip22d, Yua21]. Regularized [CLSZ21]. Regulated [LDY23]. Related [Ahm20d, Lip21h]. Reliability [Ano21a, LBTM22, LP20, Nel21, WF21]. Reliable [XQWL22]. Renewal [XKSK21]. Repairable [LP20, YCTY24]. Repeated [WM21]. Report [Apl20, Jos23]. Reproducing [LYW^{+}23]. Research [Ahm21e, Kha23, Lip21c, SA23c, Lip21g, Mar20]. Researchers [Lu20]. Resolution [MCD^{+}22, SWL^{+}22]. Response [Lip21f, Lip21e, Lip21d, NG20]. Responses [CKJD23]. Results [RRH22]. Rethinking [Mal21]. Review [Ahm20a, Ahm20b, Ahm20c, Ahm20d, Ahm21b, Ahm21a, Ahm21c, Ahm21d, Ahm21e, Ano21a, Ano21b, Ano21c, Ano21d, Ano21e, AL22, APF24, Asz23, Cha22, Che22, DM20, DSB23, Els20, Fir23, GC23, Gho20, HL23, Hus23, Kha23, Lip20a, Lip20b, Lip20d, Lip20c, Lip20e, Lip20h, Lip20f, Lip20g, Lip20i, Lip20j, Lip21b, Lip21c, Lip21g, Lip21f, Lip21e, Lip21d, Lip21h, Lip21i, Lip21j, Lip21k, Lip21l, Lip21m, Lip22c, Lip22d, Lip22a, Lip22b, Lip22e, Lip22l, Lip22f, Lip22g, Lip22h, Lip22i, Lip22k, Lip22j, Lip22m, Lip22n, Lip23a, Lip23b, Lip23f, Lip23c, Lip23d, Lip23e, Lip24a, Lip24b, Liu21, Liu23, Liu24, Lu20, Mal22a, Mal22b, Mal22c, Mal23a, Mal23b, Mal23c, Mal23d, Mal23e, MD24, Mal20, Mal21, Mar20, Mye21, Ndo24, Nel21, Oli20]. Review [Oli22, Rah23, Sa’23a, Sau21, Sau23, Sea20, SA23c, VL21a, VL21b, Wlu21, Wlu23, Zel22, ZGAR24]. Reviews [Che21, LRWD23]. Richard [Ahm20b, Mal21]. Ridge [CSX^{+}21, FL21, Has21, Hoe21, Jos21, LLBL21, Wu21, Yua21, Zou21]. Rigorous [Mal23c]. Risk [YQ20]. Ritschard [Ahm20d]. Rizzo [Lip23f]. Robert [Mal23a, Mye21]. Robin [Lip20g]. Robust [HCL23, HLPK21, LZYL22, RR21, XQ24, YK23, ZSD24]. Robustify [LCL23a]. Robustness [GSSI22]. ROC [QXY20, ZSPT21]. Rodrigo [Liu21]. Roger [Gho20]. Role [Liu24]. Ron [Lip21c]. Rong [Sa’23a]. Rosenthal [Lip22c]. Roshan [Ano22c]. Roulette [Lip23g]. Royce [Lip22c]. RR [LRWD23]. Rub [Ndo24]. Rule [Lip21o, WXJT21]. Runtime [CKJD23]. Ruppert [Cha22]. Ryan [Lip22c].

S [AL22, Lip20f, Lip21c, Mal23a, Mar20, Ndo24, Sea20]. Safety [Liu24]. Salmaso [Lip20f]. Sample [Lip20f, WF21]. Samples [Ahm20b]. Sampling [GLP22, GYZ23, VE21, YXW23, ZWX23, ZM23]. Sanchez [Lip20d, Lip20a]. Sánchez-Soriano [Lip20a]. Sandip [Lip22j]. Sarvesh [MD24]. SAS [Oli20]. Sasho [Lip22i]. Satellite [GAM23]. Satisfaction [Lip20f]. Saturated [LL21]. Scala [Lip21g]. Scalable [LP23, LPCX23]. Scale [LMN^{+}22, LPCX23, MMK^{+}22]. Scatter [ZSD24]. Scattered [PTD22]. Science [AL22, Els20, Has21, Lip20b, Lip20i, Lip21c, Lip21k, Lip21o, Mar20, SA23c, ZGAR24, Lip21l, Lip23b]. Sciences [Pou21, Gho20, Lip20j, Lip22d]. Scientific [Lip22a, Lip22b, Lip23a, Mal23a, Mal23e, Lip21l]. Scientists [Sau23]. Scope [SA23c]. Score [ZBA23]. Scott [Fir23]. Screening [FL21, QXY20, SAMJ23, VGS20, WSSE21, XJR23]. Seal [HG22]. Seasonal [WZZ23]. Second [Lip22c, Nel21]. Security [DSB23]. Selection [FL21, NG20, SS23, VGS20, Wu21, XSL^{+}23, MF21]. Self [Mal22a]. Self-Teaching [Mal22a]. Semi [YK23]. Semi-Supervised [YK23]. Semiparametric [Cha22]. SenGupta [Liu23]. Sensitivity [WBG22, Yan24]. Sensor [GMLB22, LP20]. Sentiment [LRWD23]. Sentiment-Topic [LRWD23]. Sequence [Ahm20d]. Sequences [JYYJ22]. Sequential [QX22, SBD^{+}22, WXX^{+}23, WGKJ24, WSSR21, ZH21, ZHMY22]. Serial [WT23]. Serially [QLL20, QX22]. Series [LK21, Sea20, VL21b, YA21, Lip24a]. Sertac [Ndo24]. Sets [AGC^{+}21]. Seven [Lip21n]. Shady [Ndo24]. Shandelle [SA23c]. Shape [HTLS21]. Shapes [STC^{+}22]. Shapley [Lip20a]. Sharper [LCL23b]. Sheffer [Mal23d]. Short [QLL20]. Short-Range [QLL20]. Shot [RD24]. Shrinkage [Hus23]. Signal [XQWL22]. Sijtsma [Lip22k]. Silvia [Mal23e]. Simplified [Dog21]. Simulation [CPW24, GBA^{+}21]. Simulations [WXJT21]. Singapore [Gho20, Lip22a, Lip22b, Lip23a, Liu23, Mar20]. Singmaster [Lip22a, Lip22b]. Sinha [Mar20]. site [HG22]. Size [WF21]. Slavin [Mal22a]. Sliced [SCA21]. Smallwood [Che22]. Smart [Ndo24]. Smooth [SWL^{+}22]. Smoothed [Dix22]. Snow [ZGAR24]. Social [Ano21e, Gho20, Lip20j, Lip21c, Lip22d, Mar20]. Soft [WDWZ21]. Sons [Lip21b]. Soriano [Lip20a]. Sort [LC23]. Sources [GYPS21, SWL^{+}22]. Space [ABRA20, LMN^{+}22, LYW^{+}23, ZH21]. Spaces [WGKJ24]. Sparse [CMAKA23, CKJD23, DCH^{+}22, FYQ^{+}21, GLLL21, GG21, LDY23, SWL^{+}22, WV21, WDZS22, XWYL23, ZYLS21]. Sparse-Learning [DCH^{+}22]. Spatial [LJK20, LH22, LP23, Lip22c, Mye21, QSK21, QBN21, ZWX23, VL21a]. Spatial-Temporal [Mye21]. Spatially [LK21]. Spatio [LY23, WWYC24]. Spatio-Temporal [LY23, WWYC24]. Spatiotemporal [XYWL21]. Speak [Lip20c]. Speak-Using [Lip20c]. Special [Pra22]. Spectral [LZYL22, PHRD22]. Spherical [PHRD22]. Split [CLSZ21, MGG20, JV22]. Split-Plot [MGG20]. Splitting [JV22]. Sport [ZGAR24]. Sports [Ahm21b, ZGAR24, Ahm21b, Fir23]. Spreitzer [Mal23a]. Springer [Ahm20a, Ahm20b, Ahm20d, Ahm21c, Ahm21e, Cha22, DM20, Gho20, Lip20d, Lip20c, Lip20e, Lip20h, Lip20f, Lip20j, Lip21l, Lip22h, Lip24a, Lip24b, Liu21, Liu23, Mal22c, Mal20, Mar20, Sa’23a]. Squares [WV21]. Stainless [OKF22]. Stan [Mal21]. Starter [Sea20]. Stata [Ahm21a]. State [ABRA20]. Static [LP20]. Station [WWYC24]. Statistical [APF24, Gho21b, Hus23, LZC21, LX23, Lip21e, Lip23b, Liu21, LLK^{+}23, MMO24, OKF22, PL21, QX22, STC^{+}22, SYH21, YCTY24, Zel22, ZdC21, Ahm21d, Ahm21e, HL23, Lip20i, Lip22n, Lu20, Mal22c, Mal21, Mar20, Nel21, Oli20]. Statistically [VE21]. Statistics [Ahm21b, Gho20, Gho21a, Lip20f, Lip21b, Lip21n, Lip21p, Liu23, Mal22b, Sau23, Ano21d, Ano21e, Mal20]. Steel [OKF22]. Stefano [Ano21e]. STEM [Lip21a]. Stepwise [SBD^{+}22]. Steve [Mal22a]. Steven [Lip23a]. Stewart [Mal22b]. Still [Zou21]. Stochastic [MWG23, PHRD22, ZY23, Oli22]. Story [Lip22h]. Strategies [Hus23, WSSE21]. Strategy [GLP22]. Stream [SWL^{+}22]. Streaming [GLP22, XWYL23]. Streams [LXTP20, XQWL22, ZWX23]. Stroud [Ano21b]. Structural [XWYL23]. Structure [DHJH22]. Student [Lip23b, PC20]. Student- [PC20]. Students [SA23c, SA23c]. Studer [Ahm20d]. Stuhlpfarrer [Mal23a]. Subjective [Ano21e]. Subsampling [DSW23]. Subspace [LDY23, XWYL23, ZYLS21]. Success [Ano21a]. Sufficient [WY21]. Sums [Lip24a]. Super [SWL^{+}22]. Supersaturated [JLM^{+}21, SS23, WSSE21]. Supervised [YK23, ZBA23]. Support [CZ22, GMLB22]. Surface [BA22, NG20, OKF22, ZdC21]. Surrogate [CPW24, WBG22]. Surrogates [HG22, Pou21, SGH23, ZCG21]. Survey [Ahm20b]. Surveys [Lip20f]. Survival [YA21]. Susanne [Lip20g]. Switching [CMAKA23]. Switzerland [Ahm20a, Ahm20b, Ahm20d, Ahm21c, Ahm21e, Lip20d, Lip20c, Lip20e, Lip20h, Lip20f, Lip20j, Lip21l, Lip22h, Lip24a, Lip24b, Mal22c, Mal20]. Syed [Hus23]. Symmetric [GG21]. Symplectic [Yan24]. System [Ano21c, LP20, MMK^{+}22]. Systems [LLK^{+}23, Mal23c, YCTY24]. Szekely [Lip23f].

T [Ano21c, AL22, Lip22c]. Table [Lip23g]. Tai [Asz23]. Tale [Mal23e]. Tanaka [Mal22c]. Tanaka-Ishii [Mal22c]. Taylor [Ahm21d, Ano21a, Ano21c, Ano21e, Els20, GC23, Lip20a, Lip20b, Lip20g, Lip20i, Lip21c, Lip21g, Lip21f, Lip21e, Lip21d, Lip21h, Lip21i, Lip21j, Lip21k, Lip21m, Lip22c, Lip22d, Lip22e, Lip22l, Lip22f, Lip22g, Lip22i, Lip22k, Lip22j, Lip22m, Lip22n, Lip23f, Lip23c, Lip23d, Lip23g, Mal22b, Mal21, Mye21, Oli20, Sea20, Wlu21]. Taylor [Asz23]. Teaching [Mal22a]. Techniques [Ahm21d, Lip22n]. Technologies [Ndo24]. Technology [MD24, ZGAR24]. Technometrics} [Ano21f, Ano20a, Ano22a, Ano24, Apl20]. Teik [Sa’23a]. Telecommunication [WWYC24]. Template [HPR^{+}21]. Templates [DPJS21]. Temporal [GMLB22, LY23, Mye21, PTD23, WWYC24]. Tensor [CHH^{+}23, GYPS21, GG21, HCL23, SWL^{+}22, YPLS20, ZPS23]. Tensor-on-Tensor [GYPS21]. Tensor-on-Vector [GG21]. Teoh [Sa’23a]. Term [SGG^{+}22]. Terrorism [Lip21n]. Tessellations [PGB^{+}21]. Testing [Lip21p, TCXM23, WF21]. Tests [KL20, SYH21, WM21, ZK22]. Text [DM20]. Textbook [SA23c]. Thaller [Mal23a]. Theory [Gho20, Lip21f, Lip21e, Lip21d, Lip24b, RK23, Ano21c, Mal23d]. Thinking [Lip20c]. Thomas [DM20]. Thompson [GYZ23]. Three [CMJZ21, Mal23a]. Three-Dimensional [CMJZ21, Mal23a]. Thresholds [WV23]. Throughput [LZYL22]. Thyagarajan [Ano21c]. Tidyverse [Lip20i].

Tijms [Lip23a]. Time [KL20, LK21, LX23, WDWZ21, XYWL21, XJR23, YA21, ZM23, Sea20, VL21b]. Time-Course [WDWZ21]. Times [WV23, Mal23e]. Tissue [WDWZ21]. Tissues [CMJZ21]. Toe [Sa’23a]. Tomasz [Lip22f]. Tool [Ahm20a, LCL23b]. Tools [Ahm20b, Lip21c, Lip21e]. Topic [LRWD23]. Topics [Ahm21e]. Topological [TCXM23, ZMXX23]. Tracing [Lip20j]. Traffic [XYWL21]. Transformation [LJ21]. Transition [LJ22]. Transparent [QX22]. Transport [GBA^{+}21]. Trend [KL20]. Trevor [Lip23b]. Trials [APF24, Zel22]. Tribute [Liu23]. Trimmed [WV21]. Truncation [WXJT21]. Tsai [Asz23]. TSEC [MZHW22]. Tucker [Sea20]. Tuzzi [Lip20j]. Tweedie [FYQ^{+}21, LYW^{+}23]. Two [LZC21, SS23, Tsa23, Yan24]. Two-Dimensional [LZC21]. Two-Level [SS23, Tsa23]. Two-Parameter [Yan24].

UAV [LLK^{+}23]. Uckelmann [Lip20h]. UK [Ano21b, Lip23b]. Uncertainties [HSSP22]. Uncertainty [GLLL21, GSSI22, MMK^{+}22, SBD^{+}22]. Undergraduate [SA23c]. Understand [Ahm21b]. Understanding [Lip21p, Lip21m]. Unified [FYQ^{+}21, Lip21o]. Uniform [VE21]. Universals [Mal22c]. University [Fir23, Lip23b, Mal23d, Rah23]. Urdinez [Lip21k]. USA [Mal22a]. USD [Che22]. Use [Fir23]. Useful [Ahm20a]. Using [Ahm21b, Ahm21a, Dog21, HSSP22, LJK20, Lip20c, Lip21n, Lip22e, Lip22l, NPM20, PGB^{+}21, PTD22, QBN21, SWL^{+}22, YQ21, Yan24, ZPS23, MWG23, TCXM23, ZMXX23].

V [APF24, MD24, Zel22]. V. [Ano22c]. VA [Ahm20c, Ano21d]. Valean [Lip24a]. Validation [DZ21]. Valliant [Ahm20b]. Value [DM20, Lip20a]. Valued [LCL23a]. Variable [FL21, Wu21, ZTCA21]. Variables [CSX^{+}21]. Variance [JB23, MGG20]. Variation [BA22]. Variogram [QSK21]. Varying [WDWZ21]. Varying-Coefficient [WDWZ21]. Vaughan [Lip22m]. Vector [CMAKA23, GG21]. Vectors [ZBA23]. Veda [Mal23e]. Vedic [Mal23e]. vi [Lip20c, Lip20h]. Via [YFL22, DPJS21, FL21, GR21, HG22, LXTP20, LC23, Lip20i, MF21, QSK21, RK23, SA23b, XWYL23, ZYLS21, ZBA23]. vii [DM20]. Violence [Lip20h]. Virginia [Lip20g]. Visual [Lip22i]. Visualizing [RRH22]. Vito [Lip20a]. Vogl [ZGAR24]. Voices [Ahm21e]. Vol [Lip22a, Lip22b]. Volume [Lip21f, Lip21e]. Voronoi [PGB^{+}21]. vs [Mal22c].

W [DSB23, Lip23f, Lip23c, Lip23g, Lip21i, Lip21j, Lip21k, Lip22d, Lip22e, Lip22h, Lip22k, Lip24b, Liu23, Liu24]. Wand [Cha22]. Wang [HL23]. Warfare [Lip20h]. Warjurkar [MD24]. Warranty [WZZ23]. Water [AX20]. Wayne [Fir23]. Weighted [AX20, DCW20]. Weighting [Ahm20b]. Welham [Lip22l]. Well [Ano21e]. Well-Being [Ano21e]. Wheater [Mal22a]. White [Lip23e]. Wiley [Che22, Lip21b, Lu20, Ndo24, Nel21]. Wiley-IEEE [Che22, Ndo24]. William [Ahm20c, Ano21a, Ano21d, Lu20, Nel21].

Williams [Lip22m]. Williamsburg [Ahm20c, Ano21d]. Wim [Lip21f, Lip21e, Lip21d]. Wind [AS21, Els20, PTD23]. Winston [Fir23]. Within [YXW23]. Within-and-between [YXW23]. World [Lip22a, Lip22b, Lip23a, Mal23a, Mal23e, Mal23a]. WRF [KK21].

x [Ahm20b, Ahm20d, Lip20f, Oli22]. xi [Cha22, Lip23g, Mal22b, Mal20, Mar20]. Xiali [DSB23]. xii [Ahm20a, Ano21e, Lip20j, Rah23, Wlu21]. xiii [Asz23, Lip20e, Lip22f, Nel21]. xiv [Lip22h, Lip23c]. XIX [Liu23, Lip21k, Lip23b]. Xuefeng [HL23]. xv [Gho20, Lip20g, Liu21, Oli20, Sau23, VL21b]. xvi [Lip22g, Mal21]. xvii [Lip21b, Lip23f, Lip24b, Sea20]. xviii [Lip22m]. xx1 [Els20]. XXI [SA23c, Lip22l]. xxix [Lip20a]. xxv [Lip21l]. xxvi [Ano21b, Mye21]. xxx [Lip20b, Lip20i]. xxxiii [Ahm21d, Lip22e, Lip22n]. xxxv [Lu20]. xxxviii [Lip24a].

Years [Jos21, Zou21]. Yi [Lip23c]. Yichuan [Ahm21c, Ahm21e]. Ying [APF24, Zel22]. York [APF24, DM20, HL23, Hus23, Kha23, Nel21, ZGAR24, Liu24]. You’ll [Mal22a]. Yu [Els20, Mal20]. Yüzbası [Hus23].

Zhang [Ahm21e]. Zhao [Ahm21c]. Zheng [Sa’23a]. Ziegel [Ano20b, Ano22b]. Zoccai [Ahm20a]. Zoe [Lip22l].

References

Abdalla:2020:BSS

[ABRA20] Nada Abdalla, Sudipto Banerjee, Gurumurthy Ramachandran, and Susan Arnold. Bayesian state space modeling of physical processes in industrial hygiene. Technometrics, 62(2):147–160, 2020. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Azzimonti:2021:ADE

[AGC^{+}21] Dario Azzimonti, David Ginsbourger, Clément Chevalier, Julien Bect, and Yann Richet. Adaptive design of experiments for conservative estimation of excursion sets. Technometrics, 63(1):13– 26, 2021. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Ahmed:2020:BRD

[Ahm20a] S. Ejaz Ahmed. Book review: Diagnostic Meta-Analysis: A Useful Tool for Clinical Decision-Making, by Giuseppe Biondi-Zoccai, (Ed.). Cham, Switzerland: Springer International Publishing, 2018, xii + 319 pp., $179.99, ISBN: 978-3-319-78965-1. Technometrics, 62(2):1–289, 2020. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Ahmed:2020:BRPa

[Ahm20b] S. Ejaz Ahmed. Book review: Practical Tools for Designing and Weighting Survey Samples (2nd ed.), by Richard Valliant, Jill A. Dever, and Frauke Kreuter. Cham, Switzerland: Springer, 2018, x + 776 pp., $99.99, ISBN: 978-3-319-93631-4. Technometrics, 62 (1):140, 2020. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Ahmed:2020:BRPb

[Ahm20c] S. Ejaz Ahmed. Book review: Probability (2nd ed.), by Lawrence M. Leemis. Williamsburg, VA: The College of William and Mary, 2010, ix + 555 pp., ISBN: 978-0-9829174-7-3. Technometrics, 62 (2):1–290, 2020. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Ahmed:2020:BRS

[Ahm20d] S. Ejaz Ahmed. Book review: Sequence Analysis and Related Approaches: Innovative Methods and Applications, by Gilbert Ritschard and Mattias Studer. Geneva, Switzerland: Springer, 2018, x + 298 pp, $59.98, ISBN: 978-319-95419-6. Technometrics, 62(1):139–140, 2020. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Ahmed:2021:BRD

[Ahm21a] Feryaal Ahmed. Book review: Data Management Using Stata: A Practical Handbook, by Michael N. Mitchell. Technometrics, 63 (1):136–137, 2021. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Ahmed:2021:BRA

[Ahm21b] S. Ejaz Ahmed. Book review: Analytic Methods in Sports: Using Mathematics and Statistics to Understand Data from Baseball, Football, Basketball, and Other Sports, 2nd ed. Technometrics, 63 (1):145, 2021. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Ahmed:2021:BRN

[Ahm21c] S. Ejaz Ahmed. Book review: New Frontiers of Biostatistics and Bioinformatics, edited by Yichuan Zhao and Ding-Geng Chen, Switzerland, Springer Nature, pp. 457, $136.93, ISBN: 978-3-319-99388-1. Technometrics, 63(3):441–442, 2021. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Ahmed:2021:BRSa

[Ahm21d] S. Ejaz Ahmed. Book review: Statistical and Machine-Learning Data Mining: Techniques for Better Predictive Modelling and Analysis of Big Data, edited by Bruce Ratner, Boca Raton, FL: CRC Chapman & Hall, Taylor and Francis Group, 2017, pp. xxxiii + 645, $123.83, ISBN: 978-0-367-57360-7. Technometrics, 63(2): 280, 2021. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Ahmed:2021:BRSb

[Ahm21e] S. Ejaz Ahmed. Book review: Statistical Modeling in Biomedical Research: Contemporary Topics and Voices in the Field, edited by Yichuan Zhang and Ding-Geng Chen, Switzerland, Springer Nature, 483 pp., $83.99, ISBN: 978-3-030-33415-4. Technometrics, 63 (3):442, 2021. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Arai:2022:BRM

[AL22] Kohma Arai and Vyacheslav Lyubchich. Book review: Modern Data Science with R (2nd ed.) by Benjamin S. Baumer, Daniel T. Kaplan, and Nicholas J. Horton, Boca Raton, FL: Chapman and Hall/CRC Press, 2021, 650 pp., $99.95 (Hardback), ISBN: 978-0-367-19149-8. Technometrics, 64(3):429, 2022. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Anonymous:2020:TP

[Ano20a] Anonymous. The 2018 Technometrics Prizes. Technometrics, 62 (1):141–144, 2020. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Anonymous:2020:ZAA

[Ano20b] Anonymous. 2018 Ziegel Award announcement. Technometrics, 62 (1):145, 2020. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Anonymous:2021:BRAa

[Ano21a] Anonymous. Book review: Achieving Product Reliability: A Key to Business Success, by Necip Doganaksoy, William Q. Meeker, and Gerald G. Hahn, CRC Taylor & Francis Group, Boca Raton, FL, 2021, ISBN: 978-1-138-05400-4, 217 pp., $93.18. Technometrics, 63(4):562–564, 2021. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Anonymous:2021:BRAb

[Ano21b] Anonymous. Book review: Advanced Engineering Mathematics, by K. A. Stroud and Dexter J. Booth, 6th edition. Macmillan International/Red Globe Press, London, UK, 2020. ISBN 978-1-352010-25-1 paperback, xxvi + 1222 pp., $74.99 (Advanced Engineering Mathematics — K.A. Stroud, Dexter Booth — Red Globe Press). Technometrics, 63(4):566–570, 2021. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Anonymous:2021:BRL

[Ano21c] Anonymous. Book review: Linear and Non-Linear System Theory, by T. Thyagarajan and D. Kalpana, CRC Press, Taylor & Francis Group, 2021, ISBN: 978-0-367-34014-8, 2021, 405 pp. £112.00. Technometrics, 63(4):565–566, 2021. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Anonymous:2021:BRM

[Ano21d] Anonymous. Book review: Mathematical Statistics, by Lawrence M. Leemis. The College of William and Mary Williamsburg, VA; ISBN: 978-0-982-91746-6, 2020; ix + 507 pp., $50. Technometrics, 63(4):564–565, 2021. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Anonymous:2021:BRS

[Ano21e] Anonymous. Book review: Subjective Well-Being and Social Media: Reconciling Big Data and Statistics, y Stefano M. Iacus and Giuseppe Porro. Chapman and Hall/CRC, Taylor ‘ Francis Group. Boca Raton, FL, 2021, ISBN 978-1-138-39392-9, xii + 206 pp., $99.95 (Hardback). Technometrics, 63(4):570–572, 2021. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Anonymous:2021:TEC

[Ano21f] Anonymous. Technometrics editorial collaborators. Technometrics, 63(1):146–147, 2021. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Anonymous:2022:TP

[Ano22a] Anonymous. The 2021 Technometrics Prizes. Technometrics, 64 (4):582–585, 2022. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Anonymous:2022:ZAA

[Ano22b] Anonymous. 2022 Ziegel Award announcement. Technometrics, 64 (4):586, 2022. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Anonymous:2022:EAV

[Ano22c] Anonymous. Editorial announcement, V. Roshan Joseph, Editor. Technometrics, 64(1):1, 2022. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Anonymous:2024:TAE

[Ano24] Anonymous. Technometrics 2023 Associate Editors. Technometrics, 66(1):(i), 2024. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Astaria:2024:BRH

[APF24] Tri Astaria, Yoppy Wahyu Purnomo, and Fery Muhamad Firdaus. Book review: Handbook of Statistical Methods for Randomized Controlled Trials, 1st ed.Edited by KyungMann Kim, Frank Bretz, Ying Kuen K. Cheung, Lisa V. Hampson, New York: Chapman & Hall, 2023, 654 pp., £47.99 (paperback), ISBN 978-1-032-00910-0. Technometrics, 66(1):133–134, 2024. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Apley:2020:TER

[Apl20] Daniel Apley. Technometrics 2019 Editor’s report. Technometrics, 62(1):1–5, 2020. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Agarwal:2021:BFQ

[AS21] Gaurav Agarwal and Ying Sun. Bivariate functional quantile envelopes with application to radiosonde wind data. Technometrics, 63(2):199–211, 2021. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Aszani:2023:BRM

[Asz23] Aszani Aszani. Book review: Machine Learning for Knowledge Discovery with R: Methodologies for Modeling, Inference, and Prediction, Kao-Tai Tsai, Boca Raton, FL: CRC Press, Taylor Francis Group, LLC, 2022, xiii + 260 pp., $88.00, ISBN: 978-1-032-06536-6 (H). Technometrics, 65(4):607–609, 2023. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Agarwal:2020:MBC

[AX20] Amal Agarwal and Lingzhou Xue. Model-based clustering of nonparametric weighted networks with application to water pollution analysis. Technometrics, 62(2):161–172, 2020. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Bui:2022:ANP

[BA22] Anh Tuan Bui and Daniel W. Apley. Analyzing nonparametric part-to-part variation in surface point cloud data. Technometrics, 64(4):457–474, 2022. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Chung:2024:FMO

[CA24] Seokhyun Chung and Raed Al Kontar. Federated multi-output Gaussian processes. Technometrics, 66(1):90–103, 2024. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Charles:2022:BRS

[Cha22] E. Heckler Charles. Book review: Semiparametric Regression with R, by Jaroslaw Harezlak, David Ruppert, and Matt P. Wand. Springer, 2018, ISBN: 978-1-4939-8853-2, xi + 331 pp., eBook, $87.20. Technometrics, 64(1):145, 2022. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Chen:2021:BR

[Che21] Li-Pang Chen. Book reviews. Technometrics, 63(2):272–273, 2021. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Chen:2022:BRE

[Che22] Li-Pang Chen. Book review: The ESD Control Program Handbook, by Jeremy M. Smallwood. Wiley-IEEE Press, 538 pp., USD $140.00 (hbk). ISBN 978-1-118-31103-5 2020. Technometrics, 64 (1):148–149, 2022. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Chen:2023:BNT

[CHH^{+}23] Shuoli Chen, Kejun He, Shiyuan He, Yang Ni, and Raymond K. W. Wong. Bayesian nonlinear tensor regression with functional fused elastic net prior. Technometrics, 65(4):524–536, 2023. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Chu:2021:ACC

[CJXD21] Shuyu Chu, Huijing Jiang, Zhengliang Xue, and Xinwei Deng. Adaptive convex clustering of generalized linear models with application in purchase likelihood prediction. Technometrics, 63(2): 171–183, 2021. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Chen:2023:BSR

[CKJD23] Xiaoyu Chen, Xiaoning Kang, Ran Jin, and Xinwei Deng. Bayesian sparse regression for mixed multi-responses with application to runtime metrics prediction in fog manufacturing. Technometrics, 65(2):206–219, 2023. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Chen:2021:GPA

[CKL21] Jiuhai Chen, Lulu Kang, and Guang Lin. Gaussian process assisted active learning of physical laws. Technometrics, 63(3):329– 342, 2021. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Centofanti:2021:FRC

[CLM^{+}21] Fabio Centofanti, Antonio Lepore, Alessandra Menafoglio, Biagio Palumbo, and Simone Vantini. Functional regression control chart. Technometrics, 63(3):281–294, 2021. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Christidis:2021:SRR

[CLSZ21] Anthony-Alexander Christidis, Laks Lakshmanan, Ezequiel Smucler, and Ruben Zamar. Split regularized regression. Technometrics, 62(3):330–338, 2021. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Chavez-Martinez:2023:PES

[CMAKA23] Gilberto Chavez-Martinez, Ankush Agarwal, Abbas Khalili, and Syed Ejaz Ahmed. Penalized estimation of sparse Markov regimeswitching vector auto-regressive models. Technometrics, 65(4): 553–563, 2023. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Chen:2021:FFK

[CMJZ21] Jialei Chen, Simon Mak, V. Roshan Joseph, and Chuck Zhang. Function-on-function kriging, with applications to threedimensional printing of aortic tissues. Technometrics, 63(3):384– 395, 2021. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Chan:2024:CSS

[CPW24] Moses Y.-H. Chan, Matthew Plumlee, and Stefan M. Wild. Constructing a simulation surrogate with partially observed output. Technometrics, 66(1):1–13, 2024. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Choi:2021:CRR

[CSX^{+}21] Nam-Hee Choi, Kerby Shedden, Gongjun Xu, Xuefei Zhang, and Ji Zhu. Comment: Ridge regression, ranking variables and improved principal component regression. Technometrics, 62(4):451– 455, 2021. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Chen:2022:DRC

[CZ22] Yunlu Chen and Nan Zhang. Density regression with conditional support points. Technometrics, 64(3):396–408, 2022. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Du:2022:NSL

[DCH^{+}22] Juan Du, Shanshan Cao, Jeffrey H. Hunt, Xiaoming Huo, and Jianjun Shi. A new sparse-learning model for maximum gap reduction of composite fuselage assembly. Technometrics, 64(3): 409–418, 2022. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Dong:2020:MCD

[DCW20] Hang Dong, Nan Chen, and Kaibo Wang. Modeling and change detection for count-weighted multilayer networks. Technometrics, 62(2):184–195, 2020. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Ding:2022:FPC

[DHJH22] Fei Ding, Shiyuan He, David E. Jones, and Jianhua Z. Huang. Functional PCA with covariate-dependent mean and covariance structure. Technometrics, 64(3):335–345, 2022. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Dixon:2022:IFE

[Dix22] Matthew Dixon. Industrial forecasting with exponentially smoothed recurrent neural networks. Technometrics, 64(1):114– 124, 2022. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Dai:2020:BRP

[DM20] Fan Dai and Ranjan Maitra. Book review: Practical Text Analytics: Maximizing the Value of Text Data, by Murugan Anandarajan, Chelsey Hill, and Thomas Nolan, New York, NY: Springer Publishing Company, Incorporated, 2018, vii + 285 pp., $76.64, ISBN: 978-3-319-95663-3. Technometrics, 62(2):1–286, 2020. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Doganaksoy:2021:SFL

[Dog21] Necip Doganaksoy. A simplified formulation of likelihood ratio confidence intervals using a novel property. Technometrics, 63 (1):127–135, 2021. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Dasgupta:2021:MCD

[DPJS21] Sutanoy Dasgupta, Debdeep Pati, Ian H. Jermyn, and Anuj Srivastava. Modality-constrained density estimation via deformable templates. Technometrics, 63(4):536–547, 2021. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Dharma:2023:BRA

[DSB23] Fajar Pitarsi Dharma, Moses Laksono Singgih, and Hamdan S. Bintang. Book review: AI, Machine Learning and Deep Learning a Security Perspective, edited by Fei Hu, Xiali Hei, Boca Raton, FL: CRC Press, 2023, 346 pp., 136 B/W Illustrations, GBP 99.99 (Hardback), ISBN 978-1-032-03404-1, https://doi.org/10.1201/9781003187158. Technometrics, 65(4):

606–607, 2023. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Dai:2023:SMR

[DSW23] Wenlin Dai, Yan Song, and Dianpeng Wang. A subsampling method for regression problems based on minimum energy criterion. Technometrics, 65(2):192–205, 2023. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Datta:2021:NCV

[DZ21] Abhirup Datta and Hui Zou. A note on cross-validation for lasso under measurement errors. Technometrics, 62(4):549–556, 2021. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Eck:2020:MDE

[ECNA20] Daniel J. Eck, R. Dennis Cook, Christopher J. Nachtsheim, and Thomas A. Albrecht. Multivariate design of experiments for engineering dimensional analysis. Technometrics, 62(1):6–20, 2020. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Elsayed:2020:BRD

[Els20] Elsayed A. Elsayed. Book review: Data Science for Wind Energy, by Yu Ding. Boca Raton, FL: CRC Press, Taylor & Francis Group, 2020, xx1 + 387 pp., $102.12, ISBN: 978-1-138-59052-6. Technometrics, 62(2):1–278, 2020. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Firdous:2023:BRM

[Fir23] Ahmad Mala Firdous. Book review: Mathletics: How Gamblers, Managers, and Fans Use Mathematics in Sports, 2nd ed. Wayne L. Winston, Scott Nestler, and Konstantinos Pelechrinis, Princeton University Press, 2022, 608 pp., $24.95, ISBN: 978-0-691-17762-5. Technometrics, 65(3):450–451, 2023. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Fan:2021:CFS

[FL21] Jianqing Fan and Runze Li. Comment: Feature screening and variable selection via iterative ridge regression. Technometrics, 62 (4):434–437, 2021. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Fontaine:2021:UAS

[FYQ^{+}21] Simon Fontaine, Yi Yang, Wei Qian, Yuwen Gu, and Bo Fan. A unified approach to sparse Tweedie modeling of multisource insurance claim data. Technometrics, 62(3):339–356, 2021. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Gu:2023:CIG

[GAM23] Mengyang Gu, Kyle Anderson, and Erika McPhillips. Calibration of imperfect geophysical models by multiple satellite interferograms with measurement bias. Technometrics, 65(4):453–464, 2023. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Grosskopf:2021:GCM

[GBA^{+}21] Michael Grosskopf, Derek Bingham, Marvin L. Adams, W. Daryl Hawkins, and Delia Perez-Nunez. Generalized computer model calibration for radiation transport simulation. Technometrics, 63 (1):27–39, 2021. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Garcia-Ceja:2023:BRC

[GC23] Enrique Garcia-Ceja. Book review: A Criminologist’s Guide to R: Crime by the Numbers, Jacob Kaplan, Boca Raton, FL: Chapman and Hall/CRC Press, Taylor & Francis Group, 2022, 432 pp., ISBN 978-1-032-24407-5. Technometrics, 65(4):610–611, 2023. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Guha:2021:BGS

[GG21] Sharmistha Guha and Rajarshi Guhaniyogi. Bayesian generalized sparse symmetric tensor-on-vector regression. Technometrics, 63 (2):160–170, 2021. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Gutierrez:2022:BDF

[GG22] Rene Gutierrez and Rajarshi Guhaniyogi. Bayesian dynamic feature partitioning in high-dimensional regression with big data. Technometrics, 64(2):224–240, 2022. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Ghosh:2020:BRI

[Gho20] Subir Ghosh. Book review: The Information Theory of Comparisons, With Applications to Statistics and the Social Sciences, by Roger Bowden. Singapore: Springer, 2018, xv + 159 pp., ISBN: 978-981-13-1549-7. Technometrics, 62(2):1–278, 2020. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Ghosh:2021:CIM

[Gho21a] Subir Ghosh. Computer intensive methods in statistics. Technometrics, 62(3):415, 2021. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Ghosh:2021:MAP

[Gho21b] Subir Ghosh. The measurement of association: A permutation statistical approach. Technometrics, 62(3):416, 2021. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Ghosh:2021:PBM

[Gho21c] Subir Ghosh. Probability and Bayesian modeling. Technometrics, 62(3):415–416, 2021. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Gao:2021:UQH

[GLLL21] Qi Gao, Randy C. S. Lai, Thomas C. M. Lee, and Yao Li. Uncertainty quantification for high-dimensional sparse nonparametric additive models. Technometrics, 62(4):513–524, 2021. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Gomez:2022:ASS

[GLP22] Ana María Estrada Gómez, Dan Li, and Kamran Paynabar. An adaptive sampling strategy for online monitoring and diagnosis of high-dimensional streaming data. Technometrics, 64(2):253– 269, 2022. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Grosskopf:2022:TCF

[GMLB22] Michael J. Grosskopf, Kary Myers, Earl Lawrence, and Derek Bingham. Temporal characterization and filtering of sensor data to support anomaly detection. Technometrics, 64(4):475–486, 2022. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

George:2021:CRB

[GR21] Edward I. George and Veronika Rocková. Comment: Regularization via Bayesian penalty mixing. Technometrics, 62(4):438– 442, 2021. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Gauchy:2022:IGA

[GSSI22] Clement Gauchy, Jerome Stenger, Roman Sueur, and Bertrand Iooss. An information geometry approach to robustness analysis for the uncertainty quantification of computer codes. Technometrics, 64(1):80–91, 2022. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Gahrooei:2021:MTT

[GYPS21] Mostafa Reisi Gahrooei, Hao Yan, Kamran Paynabar, and Jianjun Shi. Multiple tensor-on-tensor regression: An approach for modeling processes with heterogeneous sources of data. Technometrics, 63(2):147–159, 2021. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Guo:2023:BPO

[GYZ23] Jie Guo, Hao Yan, and Chen Zhang. A Bayesian partially observable online change detection approach with Thompson sampling. Technometrics, 65(2):179–191, 2023. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Hastie:2021:RRE

[Has21] Trevor Hastie. Ridge regularization: An essential concept in data science. Technometrics, 62(4):426–433, 2021. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Hung:2021:RFF

[HC21] Tzu-Hsiang Hung and Peter Chien. A random Fourier feature method for emulating computer models with gradient information. Technometrics, 63(4):500–509, 2021. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Heng:2023:RLR

[HCL23] Qiang Heng, Eric C. Chi, and Yufeng Liu. Robust low-rank tensor decomposition with the L_{2} criterion. Technometrics, 65(4):537– 552, 2023. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Huang:2022:MOC

[HG22] Jiangeng Huang and Robert B. Gramacy. Multi-output calibration of a honeycomb seal via on-site surrogates. Technometrics, 64(4): 548–563, 2022. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Haq:2023:BRS

[HL23] Irvanal Haq and Nila Lestari. Book review: Statistical Genomics, Brooke Fridley and Xuefeng Wang,New York, NY: Humana, 2023, 377 pp., EUR 169.99, ISBN 978-1-0716-2986-4 (eBook). Technometrics, 65(4):609–610, 2023. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Hullait:2021:RFF

[HLPK21] Harjit Hullait, David S. Leslie, Nicos G. Pavlidis, and Steve King. Robust function-on-function regression. Technometrics, 63(3): 396–409, 2021. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Hoerl:2021:RRH

[Hoe21] Roger W. Hoerl. Ridge regression: A historical context. Technometrics, 62(4):420–425, 2021. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Horton:2021:TPB

[HPR^{+}21] W. Zachary Horton, Garritt L. Page, C. Shane Reese, Lindsey K. Lepley, and McKenzie White. Template priors in Bayesian curve registration. Technometrics, 63(4):487–499, 2021. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Haizler:2021:FDO

[HS21] Tamar Haizler and David M. Steinberg. Factorial designs for online experiments. Technometrics, 63(1):1–12, 2021. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Horiguchi:2022:UBP

[HSSP22] Akira Horiguchi, Thomas J. Santner, Ying Sun, and Matthew T. Pratola. Using BART to perform Pareto optimization and quantify its uncertainties. Technometrics, 64(4):564–574, 2022. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Hu:2020:MLD

[HSZK20] Wei Hu, Weining Shen, Hua Zhou, and Dehan Kong. Matrix linear discriminant analysis. Technometrics, 62(2):196–205, 2020. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Harris:2021:EDD

[HTLS21] Trevor Harris, J. Derek Tucker, Bo Li, and Lyndsay Shand. Elastic depths for detecting shape anomalies in functional data. Technometrics, 63(4):466–476, 2021. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Hussein:2023:BRP

[Hus23] Abdulkadir Hussein. Book review: Post-Shrinkage Strategies in Statistical and Machine Learning for High Dimensional Data, Syed Ejaz Ahmed, Feryaal Ahmed, and Bahadir Yüzbaşı, New York: Chapman and Hall/CRC Press, 2023, 408 pp., ISBN 978-0-367-76344-2. Technometrics, 65(4):611, 2023. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Jin:2023:TOV

[JB23] Ying Jin and Shan Ba. Toward optimal variance reduction in online controlled experiments. Technometrics, 65(2):231–242, 2023. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Jones:2021:CPA

[JLM^{+}21] Bradley Jones, Ryan Lekivetz, Dibyen Majumdar, Christopher J. Nachtsheim, and Jonathan W. Stallrich. Construction, properties, and analysis of group-orthogonal supersaturated designs. Technometrics, 62(3):403–414, 2021. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Joseph:2021:ECY

[Jos21] V. Roshan Joseph. Editorial: Celebrating 50 years of ridge regression. Technometrics, 62(4):419, 2021. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Joseph:2023:ER

[Jos23] V. Roshan Joseph. Editor’s report. Technometrics, 65(1):1–3, 2023. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Joseph:2022:SOM

[JV22] V. Roshan Joseph and Akhil Vakayil. SPlit: An optimal method for data splitting. Technometrics, 64(2):166–176, 2022. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Jin:2022:BHM

[JYYJ22] Huaqing Jin, Guosheng Yin, Binhang Yuan, and Fei Jiang. Bayesian hierarchical model for change point detection in multivariate sequences. Technometrics, 64(2):177–186, 2022. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Khan:2023:BRE

[Kha23] Nisar Ahmad Khan. Book review: The Effect: An Introduction to Research Design and Causality, by Nisar Ahmad Khan, New York, Chapman and Hall/CRC, 2021, 646 pp., $23.99 (pbk), ISBN 978-1-032-12578-7. Technometrics, 65(2):297–298, 2023. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Konomi:2021:BAM

[KK21] Bledar A. Konomi and Georgios Karagiannis. Bayesian analysis of multifidelity computer models with local features and nonnested experimental designs: Application to the WRF model. Technometrics, 63(4):510–522, 2021. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Kvaloy:2020:CTT

[KL20] Jan Terje Kvaløy and Bo Henry Lindqvist. A class of tests for trend in time censored recurrent event data. Technometrics, 62 (1):101–115, 2020. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Khakifirooz:2021:MMG

[KTF21] Marzieh Khakifirooz, Sheng Tsaing Tseng, and Mahdi Fathi. Model misspecification of generalized gamma distribution for accelerated lifetime-censored data. Technometrics, 62(3):357–370, 2021. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Lewis-Beck:2022:PFF

[LBTM22] Colin Lewis-Beck, Qinglong Tian, and William Q. Meeker. Prediction of future failures for heterogeneous reliability field data.

Technometrics, 64(1):125–138, 2022. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Lin:2023:AOA

[LC23] Dennis K. J. Lin and Jianbin Chen. Adaptive order-of-addition experiments via the quick-sort algorithm. Technometrics, 65(3): 396–405, 2023. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Lei:2023:FOD

[LCL23a] Xinyi Lei, Zhicheng Chen, and Hui Li. Functional outlier detection for density-valued data with application to robustify distributionto-distribution regression. Technometrics, 65(3):351–362, 2023. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Liu:2023:SCT

[LCL23b] Xiaoqian Liu, Eric C. Chi, and Kenneth Lange. A sharper computational tool for regression. Technometrics, 65(1):117–126, 2023. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Li:2023:DAC

[LCX23] Chunya Li, Xiaojun Cui, and Shifeng Xiong. Design and analysis of computer experiments with both numeral and distributional inputs. Technometrics, 65(3):406–417, 2023. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Liu:2023:CRS

[LDY23] Xingchen Liu, Juan Du, and Zhi-Sheng Ye. A covariate-regulated sparse subspace learning model and its application to process monitoring and fault isolation. Technometrics, 65(2):269–280, 2023. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Lee:2022:PEE

[LH22] Ben Seiyon Lee and Murali Haran. PICAR: An efficient extendable approach for fitting hierarchical spatial models. Technometrics, 64 (2):187–198, 2022. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Lipovetsky:2020:BRH

[Lip20a] Stan Lipovetsky. Book review: Handbook of the Shapley Value, by Encarnación Algaba, Vito Fragnelli, and Joaquín Sánchez-Soriano, editors. Boca Raton, FL: Chapman and Hall/CRC, Taylor & Francis Group, 2020, xxix + 576 pp., $160.00, ISBN: 978-0-8153-7468-8. Technometrics, 62(2):1–280, 2020. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Lipovetsky:2020:BRI

[Lip20b] Stan Lipovetsky. Book review: Introduction to Data Science: Data Analysis and Prediction Algorithms With R, by Rafael A. Irizarry. Boca Raton, FL: Chapman and Hall/CRC, Taylor & Francis Group, 2020, xxx + 713 pp., $99.95, ISBN: 978-0-367-35798-6. Technometrics, 62(2):1–282, 2020. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Lipovetsky:2020:BRLb

[Lip20c] Stan Lipovetsky. Book review: Let the Evidence Speak-Using Bayesian Thinking in Law, Medicine, Ecology and Other Areas, by Alan Jessop. Cham, Switzerland: Springer, 2018, vi + 226 pp., $37.99, ISBN: 978-3-319-71391-5. Technometrics, 62(1):137– 138, 2020. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Lipovetsky:2020:BRLa

[Lip20d] Stan Lipovetsky. Book review: Lexical Collocation Analysis: Advances and Applications, by Pascual Cantos-Gomez and Moises Almela-Sanchez. Cham, Switzerland: Springer, 2018, ix + 140 pp., $119.99, ISBN: 978-319-92581-3. Technometrics, 62(1):137, 2020. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Lipovetsky:2020:BRM

[Lip20e] Stan Lipovetsky. Book review: Modern Psychometrics With R, by Patrick Mair. Cham, Switzerland: Springer, 2018, xiii + 458 pp., $74.99, ISBN: 978-3-319-93175-3. Technometrics, 62(1):135– 137, 2020. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Lipovetsky:2020:BRPb

[Lip20f] Stan Lipovetsky. Book review: Parametric and Nonparametric Statistics for Sample Surveys and Customer Satisfaction Data, by R. Arboretti, A. Bathke, S. Bonnini, P. Bordignon, E. Carrozzo, L. Corain, and L. Salmaso. Cham, Switzerland: Springer, 2018, x + 84 pp., $69.99, ISBN: 978-319-91739-9. Technometrics, 62 (1):138–139, 2020. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Lipovetsky:2020:BRPc

[Lip20g] Stan Lipovetsky. Book review: Practical Multivariate Analysis (6th ed.), by Abdelmonem Afifi, Susanne May, Robin Donatello, and Virginia A. Clark. Boca Raton, FL: Chapman and Hall/ CRC, Taylor & Francis Group, 2020, xv + 418 pp., $99.95, ISBN: 978-1-138-70222-6. Technometrics, 62(2):1–283, 2020. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Lipovetsky:2020:BRPa

[Lip20h] Stan Lipovetsky. Book review: Prehistoric Warfare and Violence: Quantitative and Qualitative Approaches, by Andrea Dolfini, Rachel J. Crellin, Christian Horn, and Marion Uckelmann. Cham, Switzerland: Springer, 2018, vi + 365 pp., $129.00, ISBN: 978-319-78827-2. Technometrics, 62(1):132–135, 2020. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Lipovetsky:2020:BRS

[Lip20i] Stan Lipovetsky. Book review: Statistical Inference via Data Science: A Modern Dive Into R and the Tidyverse, by Chester Ismay and Albert Y. Kim. Boca Raton, FL: Chapman and Hall/CRC, Taylor & Francis Group, 2020, xxx + 430 pp., $79.95, ISBN: 978-0-367-40982-1. Technometrics, 62(2):1–283, 2020. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Lipovetsky:2020:BRT

[Lip20j] Stan Lipovetsky. Book review: Tracing the Life Cycle of Ideas in the Humanities and Social Sciences, by Arjuna Tuzzi. Cham, Switzerland: Springer, 2018, xii + 217 pp., $89.00, ISBN: 978-319-97063-9. Technometrics, 62(1):130–132, 2020. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Lipovetsky:2021:PSE

[Lip21a] Stan Lipovetsky. 2^{5} problems for STEM education. Technometrics, 62(4):557–558, 2021. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Lipovetsky:2021:BRA

[Lip21b] Stan Lipovetsky. Book review: Advanced Statistics with Applications in R, by Eugene Demidenko, Hoboken, NJ: John Wiley & Sons, 2020, xvii + 856 pp., $135.00 (Hardback), ISBN 978-1-118-38798-6. Technometrics, 63(2):273–275, 2021. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Lipovetsky:2021:BRB

[Lip21c] Stan Lipovetsky. Book review: Big Data and Social Science: Data Science Methods and Tools for Research and Practice, 2nd ed., edited by Ian Foster, Rayid Ghani, Ron S. Jarmin, Frauke Kreuter and Julia Lane, Boca Raton, FL, Chapman and Hall/CRC, Taylor & Francis Group, 2021, 411 pp., 64 color ill., $160.00 (Hardback), ISBN: 978-03-6734-187-9. Technometrics, 63(3):421–423, 2021. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Lipovetsky:2021:BRHd

[Lip21d] Stan Lipovetsky. Book review: Handbook of Item Response Theory: Applications (3), edited by Wim J. van der Linden, Boca Raton, FL, Chapman and Hall/CRC, Taylor & Francis Group, 2019, 607 pp., $59.96 (Paperback). ISBN: 978-0-367-22118-8. Technometrics, 63(3):433–437, 2021. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Lipovetsky:2021:BRHc

[Lip21e] Stan Lipovetsky. Book review: Handbook of Item Response Theory, Statistical Tools, Volume 2, edited by Wim J. van der Linden, Boca Raton, FL, Chapman and Hall/CRC, Taylor & Francis Group, 2019, 452 pp., $59.96 (Paperback), ISBN 978-0-367-22104-1. Technometrics, 63(3):431–433, 2021. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Lipovetsky:2021:BRHb

[Lip21f] Stan Lipovetsky. Book review: Handbook of Item Response Theory, Volume 1, Models, edited by Wim J. van der Linden. Boca Raton, FL, Chapman and Hall/CRC, Taylor & Francis Group, 2019, 624 pp., $75.95 (Paperback), ISBN 978-0-367-22120-3. Technometrics, 63(3):428–431, 2021. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Lipovetsky:2021:BRHa

[Lip21g] Stan Lipovetsky. Book review: Handbook of Military and Defense Operations Research, edited by Natalie M. Scala and James P. Howard, II. Boca Raton, FL: Chapman and Hall/CRC, Taylor & Francis Group, 2020, 463 pp., $152.00 (hardback), $46.36 (eBook), ISBN: 978-1-138-60733-0. Technometrics, 62(4):559, 2021. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Lipovetsky:2021:BRI

[Lip21h] Stan Lipovetsky. Book review: Integrals Related to the Error Function, by Nikolai E. Korotkov and Alexander N. Korotkov. Boca Raton, FL: Chapman and Hall/CRC Press, Taylor & Francis Group, 2020, 228 pp., $140.00 (hardback), $46.36 (eBook), ISBN: 978-0-367-40820-6 (hardback). Technometrics, 62(4):560, 2021. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Lipovetsky:2021:BRLa

[Lip21i] Stan Lipovetsky. Book review: Learning Microeconometrics with R, by Christopher P. Adams, Chapman and Hall/CRC, Taylor & Francis Group, Boca Raton, FL, 2021, ISBN 978-0-367-25538-1, 398 pp., 72 b/w illustrations, $99.95. Technometrics, 63(3):424– 425, 2021. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Lipovetsky:2021:BRLb

[Lip21j] Stan Lipovetsky. Book review: Linear Models with Python, Faraway Julian J.. Boca Raton, FL, Chapman and Hall/CRC, Taylor & Francis Group, 2021, 308 pp., 85 b/w illustrations, $99.95 (Hardback), ISBN: 978-1-138-48395-8. Technometrics, 63(3):426– 427, 2021. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Lipovetsky:2021:BRR

[Lip21k] Stan Lipovetsky. Book review: R for Political Data Science: A Practical Guide, edited by Francisco Urdinez and Andrés Cruz, Chapman and Hall/CRC, Taylor & Francis Group, Boca Raton, FL, 2021, ISBN 978-0-367-81889-0, xix + 440 pp., 209 b/w illustrations, $99.95 (Hardback). Technometrics, 63(2):277–278, 2021. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Lipovetsky:2021:BRS

[Lip21l] Stan Lipovetsky. Book review: Scientific Journeys: A Physicist Explores the Culture, History and Personalities of Science, by H. Frederick Dylla, Switzerland AG: Springer International Publisher, Springer Nature, 2020, xxv + 222 pp., $129.00 (Paperback), ISBN 978-3-030-55799-7. Technometrics, 63(2):275–277, 2021. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Lipovetsky:2021:BRU

[Lip21m] Stan Lipovetsky. Book review: Understanding the Analytic Hierarchy Process, by Konrad Kulakowski, Boca Raton, FL: Chapman and Hall/CRC, Taylor & Francis Group, 2021, 262 pp., $130.00 (Hardback), ISBN 978-1-1380-3232-3. Technometrics, 63(2):278– 279, 2021. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Lipovetsky:2021:DST

[Lip21n] Stan Lipovetsky. Debunking seven terrorism myths using statistics. Technometrics, 63(1):137–140, 2021. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Lipovetsky:2021:EKB

[Lip21o] Stan Lipovetsky. The equation of knowledge: From Bayes’ rule to a unified philosophy of science. Technometrics, 63(1):140–143, 2021. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Lipovetsky:2021:UET

[Lip21p] Stan Lipovetsky. Understanding elections through statistics: Polling, prediction, and testing. Technometrics, 63(1):143–144, 2021. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Lipovetsky:2022:BRAc

[Lip22a] Stan Lipovetsky. Book review: Adventures in Recreational Mathematics, Vol. 1 by David Singmaster, Singapore: World Scientific, 2021, 316 pp., $38.00 (pbk), ISBN 978-981-12-2650-2. Technometrics, 64(4):575–581, 2022. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Lipovetsky:2022:BRAd

[Lip22b] Stan Lipovetsky. Book review: Adventures in Recreational Mathematics, Vol. 2, by David Singmaster, Singapore: World Scientific, 2022, 180 pp., $28.00 (pbk), ISBN 978-981-12-2651-9. Technometrics, 64(4):576–578, 2022. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Lipovetsky:2022:BRAa

[Lip22c] Stan Lipovetsky. Book review: Analyzing Spatial Models of Choice and Judgment, Second Edition, by David A. Armstrong II, Ryan Bakker, Royce Carroll, Christopher Hare, Keith T. Poole, and Howard Rosenthal. CRC Press. Taylor & Francis Group, Chapman and Hall, Boca Raton, FL, 2021, ISBN 978-1-138-71533-2, 320 pp., $63.96 (hbk). Technometrics, 64(1):139–143, 2022. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Lipovetsky:2022:BRAb

[Lip22d] Stan Lipovetsky. Book review: Applied Regularization Methods for the Social Sciences by Holmes Finch, Boca Raton, FL, CRC/Chapman & Hall, Taylor & Francis Group, 2022, 305 pp., 77 b/w ill., $89.95 (hbk), ISBN 978-0-367-40878-7. Technometrics, 64(3):419–421, 2022. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Lipovetsky:2022:BRBa

[Lip22e] Stan Lipovetsky. Book review: Behavior Analysis with Machine Learning Using R, by Enrique Garcia Ceja, Boca Raton, FL, CRC Press, Taylor & Francis Group, 2022, xxxiii + 397 pp., 96 color & 88 b/w illustrations, $79.96, ISBN 978-1-032-06704-9. Technometrics, 64(3):421–423, 2022. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Lipovetsky:2022:BRE

[Lip22f] Stan Lipovetsky. Book review: Explanatory Model Analysis: Explore, Explain and Examine Predictive Models, by Przemyslaw Biecek, Tomasz Burzykowski, Boca Raton, FL, Chapman and Hall/CRC, Taylor & Francis Group, 2021, xiii + 311 pp., $79.96 (hbk), ISBN 978-0-367-13559-1. Technometrics, 64(3):423–424, 2022. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Lipovetsky:2022:BRH

[Lip22g] Stan Lipovetsky. Book review: Handbook of Regression Modeling in People Analytics: With Examples in R and Python, by Keith McNulty. CRC Press, Taylor & Francis Group, Boca Raton, FL, 2021, ISBN 978-1-032-04174-2, xvi + 255 pp., 48 color illustrations, $63.96 (hbk). Technometrics, 64(1):143–145, 2022. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Lipovetsky:2022:BRMa

[Lip22h] Stan Lipovetsky. Book review: Manhattan Project: The Story of the Century, by Bruce Cameron Reed. Springer Nature Switzerland AG, 2020, Cham, Switzerland. ISBN 978-3-030-45733-4, xiv + 553 pp., 127 b/w and 25 color illustrations, $49.99 (hbk). Technometrics, 64(2):270–278, 2022. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Lipovetsky:2022:BRMb

[Lip22i] Stan Lipovetsky. Book review: Math and Art: an Introduction to Visual Mathematics, 2nd ed., by Sasho Kalajdzievski, Boca Raton, FL, Chapman and Hall/CRC Press, Taylor & Francis Group, 2022, 398 pp., $55.96 (pbk), ISBN 978-0-367-07611-5. Technometrics, 64(3):425–426, 2022. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Lipovetsky:2022:BRMd

[Lip22j] Stan Lipovetsky. Book review: Mathematical Modeling: Models, Analysis and Applications, 2nd Edition by Sandip Banerjee, Boka Raton, FL: Chapman and Hall/CRC Press, Taylor & Francis Group, 2022, 433 pp., $130.00 (hbk), ISBN 978-1-138-49594-4. Technometrics, 64(4):580–581, 2022. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Lipovetsky:2022:BRMc

[Lip22k] Stan Lipovetsky. Book review: Measurement Models for Psychological Attributes by Klaas Sijtsma, and Andries L. van der Ark, Boca Raton, FL, CRC Press, Taylor & Francis Group, Chapman and Hall, 2020, 428 pp., 50 b/w illustrations, $79.95 (pbk), ISBN 978-0-367-42452-7. Technometrics, 64(3):426–429, 2022. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Lipovetsky:2022:BRBb

[Lip22l] Stan Lipovetsky. Book review: Multivariate Data Integration Using R: Methods and Applications with the mixOmics Package, by Kim-Ahn Lê Cao and Zoe Welham, Boca Raton, FL, CRC/CRC Press, Chapman and Hall, Taylor & Francis Group, 2022, xxi + 308 pp., 121 color illustrations, $79.96 (hbk), ISBN 978-0-367-46094-5. Technometrics, 64(3):429–431, 2022. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Lipovetsky:2022:BRP

[Lip22m] Stan Lipovetsky. Book review: Probability, Choice, and Reason, by Leighton Vaughan Williams. CRC Press/Chapman and Hall, Taylor & Francis Group, Boca Raton, FL, 2022, ISBN 978-0-367-53891-0, xviii + 292 pp., $66.95 (pbk). Technometrics, 64(1):149– 150, 2022. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Lipovetsky:2022:BRS

[Lip22n] Stan Lipovetsky. Book review: Statistical and Machine-Learning Data Mining: Techniques for Better Predictive Modeling and Analysis of Big Data, 3rd Edition, by Bruce Ratner. Chapman and Hall/CRC, Taylor & Francis Group, Boca Raton, FL, 2020; ISBN 978-0-367-57360-7; xxxiii + 655 pp., $54.95 (pbk). Technometrics, 64(1):145–148, 2022. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Lipovetsky:2023:BRCa

[Lip23a] Stan Lipovetsky. Book review: Chance, Logic and Intuition: An Introduction to the Counter-Intuitive Logic of Chance by Steven Tijms, Singapore: World Scientific, 2021, 256 pp., $34.00 (pbk), ISBN 978-981-12-4783-5. Technometrics, 65(1):134–136, 2023. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Lipovetsky:2023:BRCb

[Lip23b] Stan Lipovetsky. Book review: Computer Age Statistical Inference: Algorithms, Evidence, and Data Science, Student ed., Bradley Efron and Trevor Hastie, UK: Cambridge University Press, 2021, xix + 491 pp., $39.99 (pbk), ISBN 978-1-108-82341-8. Technometrics, 65(4):611–613, 2023. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Lipovetsky:2023:BRH

[Lip23c] Stan Lipovetsky. Book review: Handbook of Measurement Error Models, edited by Grace Y. Yi, Aurore Delaigle, and Paul Gustafson, Boca Raton, FL, Chapman & Hall/CRC Press, Taylor & Francis Group, 2021, xiv+577 pp., 33 B/W illustrations, $200.00 (hbk), ISBN 978-1-138-10640-6. Technometrics, 65(2): 302–304, 2023. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Lipovetsky:2023:BRI

[Lip23d] Stan Lipovetsky. Book review: Introduction to Math Olympiad Problems, by Michael A. Radin, Boca Raton, FL, Chapman and Hall/CRC Press, Taylor & Francis Group, 2021, 160 pp., $23.96 (pbk), ISBN 978-0-367-54471-3. Technometrics, 65(2):296, 2023. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Lipovetsky:2023:BRL

[Lip23e] Stan Lipovetsky. Book review: Luck, Logic, and White Lies: The Mathematics of Games; 2nd ed. Technometrics, 65(4):604–606, 2023. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Lipovetsky:2023:BRE

[Lip23f] Stan Lipovetsky. Book review: The Energy of Data and Distance Correlation, Gabor J. Szekely and Maria L. Rizzo, Boca Raton, FL: CRC/CRC Press/Chapman & Hall/Taylor & Francis Group, 2023, xvii + 448 pp., $120.00 (hbk), 7 Color & 24 B/W Illustrations, ISBN 978-1-482-24274-4. Technometrics, 65(3):446– 448, 2023. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Lipovetsky:2023:MBF

[Lip23g] Stan Lipovetsky. Mathematics of The Big Four Casino Table Games: Blackjack, Baccarat, Craps, & Roulette, Mark Bollman,Boca Raton, FL: CRC Press/Chapman & Hall, Taylor & Francis Group, 2021, xi + 353 pp., 43 B/W illustrations, $31.16 (pbk), ISBN 978-0-367-74090-0. Technometrics, 65(4):613–614, 2023. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Lipovetsky:2024:BRA

[Lip24a] Stan Lipovetsky. Book review: (Almost) Impossible Integrals, Sums, and Series, Cornel Ioan Valean, Cham, Switzerland: Springer Nature Switzerland AG, 2019, xxxviii + 539 pp., $69.99 (hbk), ISBN 978-3-030-02461-1. Technometrics, 66(1):139, 2024. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Lipovetsky:2024:BRQ

[Lip24b] Stan Lipovetsky. Book review: Quantitative Investing: from Theory to Industry, Lingjie Ma, Cham, Switzerland: Springer Nature Switzerland AG, 2020, xvii + 455 pp., 35 b/w and 110 color illustrations, $89.99 (pbk), ISBN 978-3-030-47204-7. Technometrics, 66(1):134–136, 2024. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Liu:2021:BRM

[Liu21] Shin Ta Liu. Book review: Machine Learning: A Practical Approach on the Statistical Learning, by Rodrigo Fernandes de Mello and Moacir Antonelli Pontintor. Springer International Publishing AG, part of Springer Nature, 2018, xv + 362 pp., ISBN: 978-3-319-94988-8. Technometrics, 62(4):560–561, 2021. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Liu:2023:BRD

[Liu23] Shuangzhe Liu. Book review: Directional Statistics for Innovative Applications: A Bicentennial Tribute to Florence Nightingale by Ashis SenGupta and Barry C. Arnold, Singapore: Springer, 2022, XIX, 488 pp., with 49 b/w illustrations, 93 illustrations in color, ISBN 978-981-19-1043-2 (hbk). Technometrics, 65(1):133– 134, 2023. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Liu:2024:BRB

[Liu24] Shuangzhe Liu. Book review: Backseat Driver: The Role of Data in Great Car Safety Debates norma faris hubele, new York: Chapman and Hall/CRC, 2022, 200 pages 67 B/W illustrations, $88, hardcover, ISBN 978-0-367-47407-2 (hbk); 978-0-367-47230-6 (pbk); 978-1-003-03534-3 (ebk). Technometrics, 66(1):131–132, 2024. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Lin:2021:TAG

[LJ21] Li-Hsiang Lin and V. Roshan Joseph. Transformation and additivity in Gaussian processes. Technometrics, 62(4):525–535, 2021. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Lee:2022:ITL

[LJ22] Seokho Lee and Hyelim Jung. Individual transition label noise logistic regression in binary classification for incorrectly labeled data. Technometrics, 64(1):18–29, 2022. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Lee:2020:BNM

[LJK20] Youngmin Lee, Taewon Jeong, and Heeyoung Kim. A Bayesian nonparametric mixture measurement error model with application to spatial density estimation using mobile positioning data with multi-accuracy and multi-coverage. Technometrics, 62(2): 173–183, 2020. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Lee:2021:BNJ

[LK21] Youngmin Lee and Heeyoung Kim. Bayesian nonparametric joint mixture model for clustering spatially correlated time series. Technometrics, 62(3):313–329, 2021. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Landgraf:2021:GPC

[LL21] Andrew J. Landgraf and Yoonkyung Lee. Generalized principal component analysis: Projection of saturated model parameters. Technometrics, 62(4):459–472, 2021. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Le:2021:CRR

[LLBL21] Can M. Le, Keith Levin, Peter J. Bickel, and Elizaveta Levina. Comment: Ridge regression and regularization of large matrices. Technometrics, 62(4):443–446, 2021. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Liu:2023:SLN

[LLK^{+}23] Xinchao Liu, Xiao Liu, Tulin Kaman, Xiaohua Lu, and Guang Lin. Statistical learning for nonlinear dynamical systems with applications to aircraft-UAV collisions. Technometrics, 65(4):564– 578, 2023. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Lee:2022:ADL

[LMN^{+}22] Wesley Lee, Tyler H. McCormick, Joshua Neil, Cole Sodja, and Yanran Cui. Anomaly detection in large-scale networks with latent space models. Technometrics, 64(2):241–252, 2022. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Liu:2020:ALH

[LP20] Xiao Liu and Rong Pan. Analysis of large heterogeneous repairable system reliability data with static system attributes and dynamic sensor measurement in big data environment. Technometrics, 62 (2):206–222, 2020. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Lee:2023:SPA

[LP23] Benjamin Seiyon Lee and Jaewoo Park. A scalable partitioned approach to model massive nonstationary non-Gaussian spatial datasets. Technometrics, 65(1):105–116, 2023. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Li:2023:SGP

[LPCX23] Yongxiang Li, Yuting Pu, Changming Cheng, and Qian Xiao. A scalable Gaussian process for large-scale periodic data. Technometrics, 65(3):363–374, 2023. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Liang:2023:JRM

[LRWD23] Qiao Liang, Shyam Ranganathan, Kaibo Wang, and Xinwei Deng. JST-RR model: Joint modeling of ratings and reviews in sentiment-topic prediction. Technometrics, 65(1):57–69, 2023. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Li:2022:GPE

[LT22] Zhaohui Li and Matthias Hwai Yong Tan. A Gaussian process emulator based approach for Bayesian calibration of a functional input. Technometrics, 64(3):299–311, 2022. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Lu:2020:BRS

[Lu20] Lu Lu. Book review: Statistical Intervals: A Guide for Practitioners and Researchers (2nd ed.), by William Q. Meeker, Gerald J. Hahn, and Luis A. Escobar. Hoboken, NJ: Wiley, 2017, xxxv + 579 pp., $87.22, ISBN: 978-0-471-68717-7. Technometrics, 62(2): 1–285, 2020. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Lu:2021:GPM

[LWHY21] Lu Lu, Bingxing Wang, Yili Hong, and Zhisheng Ye. General path models for degradation data with multiple characteristics and covariates. Technometrics, 63(3):354–369, 2021. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Li:2023:CTS

[LX23] Chenglong Li and Xun Xiao. On censoring time in statistical monitoring of lifetime data. Technometrics, 65(3):418–431, 2023. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Li:2020:DPH

[LXTP20] Wendong Li, Dongdong Xiang, Fugee Tsung, and Xiaolong Pu. A diagnostic procedure for high-dimensional data streams via missed discovery rate control. Technometrics, 62(1):84–100, 2020. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Liu:2023:IME

[LY23] Xiao Liu and Kyongmin Yeo. Inverse models for estimating the initial condition of spatio-temporal advection-diffusion processes. Technometrics, 65(3):432–445, 2023. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Lyubchich:2021:ECM

[Lyu21] Vyacheslav Lyubchich. Elements of copula modeling with R. Technometrics, 62(3):416, 2021. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Lian:2023:TCP

[LYW^{+}23] Yi Lian, Archer Yi Yang, Boxiang Wang, Peng Shi, and Robert William Platt. A Tweedie compound Poisson model in reproducing kernel Hilbert space. Technometrics, 65(2):281–295, 2023. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Lee:2021:SMA

[LZC21] Jaesung Lee, Shiyu Zhou, and Junhong Chen. Statistical modeling and analysis of k-layer coverage of two-dimensional materials in inkjet printing processes. Technometrics, 63(3):410–420, 2021. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Lysy:2022:REP

[LZYL22] Martin Lysy, Feiyu Zhu, Bryan Yates, and Aleksander Labuda. Robust and efficient parametric spectral density estimation for high-throughput data. Technometrics, 64(1):30–51, 2022. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Mallik:2020:BRP

[Mal20] Abhirup Mallik. Book review: A Parametric Approach to a Nonparametric Statistics, by Mayer Alvo and Philip L. H. Yu. Cham, Switzerland: Springer, 2018, xi + 217 pp., $50.78, ISBN: 978-3-319-94152-3. Technometrics, 62(2):1–287, 2020. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Mallik:2021:BRS

[Mal21] Abhirup Mallik. Book review: Statistical Rethinking: A Bayesian Course with Examples in R and Stan, edited by Richard McElreath, Boca Raton, FL: CRC Taylor and Francis, 2020, xvi + 589 pp., $94.94 (Hardback), ISBN: 978-0-367-13991-9. Technometrics, 63(3):440–441, 2021. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Mala:2022:BRA

[Mal22a] Firdous Ahmad Mala. Book review: All the Math You’ll Ever Need: a Self-Teaching Guide (3rd Edition) by Carolyn C. Wheater and Steve Slavin, USA: Jossey-Bass, 2022, 336 pp., $25.00 (PBK), ISBN: 978-1-119-71918-2. Technometrics, 64(4):579–580, 2022. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Mala:2022:BRB

[Mal22b] Firdous Ahmad Mala. Book review: Basic Statistics and Epidemiology: a Practical Guide (5th Edition) by Antony Stewart, CRC Press, Taylor and Francis Group, 2022, xi + 204 pp., $99.00 (HB), ISBN: 978-0-367-70815-3. Technometrics, 64(4):578–579, 2022. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Mala:2022:BRS

[Mal22c] Firdous Ahmad Mala. Book review: Statistical Universals of Language: Mathematical Chance vs. Human Choice by Kumiko Tanaka-Ishii, Springer, Switzerland, 2021, 236 pp., EUR 89.99 (hardback), EUR 76.99 (eBook), ISBN: 978-3-030-59376-6. Technometrics, 64(3):432–433, 2022. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Mala:2023:BRG

[Mal23a] Firdous Ahmad Mala. Book review: Geometry in Our Three-Dimensional World by Alfred S. Posamentier, Bernd Thaller, Christian Dorner, Robert Geretschläger, Guenter Maresch, Christian Spreitzer, David Stuhlpfarrer, World Scientific, 2021, 440 pp., $118.00 (HB), ISBN 978-981-12-3710-2. Technometrics, 65 (1):127–129, 2023. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Mala:2023:BRM

[Mal23b] Firdous Ahmad Mala. Book review: Mathematical Analysis and Optimization for Economists, by Michael J. Panik, Chapman & Hall, 2021, 345 pp., $52.49 (HB), ISBN: 978-0-367-75901-8. Technometrics, 65(2):300–301, 2023. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Mala:2023:BRN

[Mal23c] Firdous Ahmad Mala. Book review: Number Systems: A Path into Rigorous Mathematics, by Anthony Kay, CRC Press, 2022, 317 pp., $130.00 (hcb), ISBN: 978-0-367-18061-4. Technometrics, 65(2):301–302, 2023. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Mala:2023:BRP

[Mal23d] Firdous Ahmad Mala. Book review: Polynomial Methods and Incidence Theory, Adam Sheffer, Cambridge University Press, 2022, 263 pp., $64.99, ISBN: 978-1-108-95998-8. Technometrics, 65(3): 449–450, 2023. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Mala:2023:BRV

[Mal23e] Firdous Ahmad Mala. Book review: Vedic Mathematics: A Mathematical Tale from the Ancient Veda to Modern Times by Giuseppe Dattoli, Silvia Licciardi, and Marcello Artioli, World Scientific, 2021, 232 pp., $68 (HB), ISBN 978-981-12-2155-2. Technometrics, 65(1):131–132, 2023. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Marzjarani:2020:BRS

[Mar20] Morteza Marzjarani. Book review: Statistical Methods in Social Science Research, by S. P. Mukherjee, Bikas K. Sinha, and Asis Chatterjee. Singapore: Springer Nature, 2018, xi + 152 pp., $95.41, ISBN: 978-981-13-2145-0, ISBN: 978-981-13-2146-7 (ebook). Technometrics, 62(2):1–288, 2020. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Miller:2022:ALM

[MCD^{+}22] Matthew J. Miller, Matthew J. Cabral, Elizabeth C. Dickey, James M. LeBeau, and Brian J. Reich. Accounting for location measurement error in imaging data with application to atomic resolution images of crystalline materials. Technometrics, 64(1): 103–113, 2022. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Mala:2024:BRE

[MD24] Firdous Ahmad Mala and Shahid Abdullah Dar. Book review: Ethics in Information Technology: a Practical Guide G. K. Awari and Sarvesh V. Warjurkar, Boca Raton, FL: CRC Press, 2022, 213 pp., $53.59, ISBN 978-1-032-16379-6. Technometrics, 66(1):132– 133, 2024. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Metzger:2021:DLH

[MF21] Thomas A. Metzger and Christopher T. Franck. Detection of latent heteroscedasticity and group-based regression effects in linear models via Bayesian model selection. Technometrics, 63(1):116– 126, 2021. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Mylona:2020:OBS

[MGG20] Kalliopi Mylona, Steven G. Gilmour, and Peter Goos. Optimal blocked and split-plot designs ensuring precise pure-error estimation of the variance components. Technometrics, 62(1):57–70, 2020. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Ma:2022:CME

[MMK^{+}22] Pulong Ma, Anirban Mondal, Bledar A. Konomi, Jonathan Hobbs, Joon Jin Song, and Emily L. Kang. Computer model emulation with high-dimensional functional output in large-scale observing system uncertainty experiments. Technometrics, 64(1): 65–79, 2022. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Malinovskaya:2024:SPM

[MMO24] Anna Malinovskaya, Pavlo Mozharovskyi, and Philipp Otto. Statistical process monitoring of artificial neural networks. Technometrics, 66(1):104–117, 2024. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Munezero:2023:DME

[MVK23] Parfait Munezero, Mattias Villani, and Robert Kohn. Dynamic mixture of experts models for online prediction. Technometrics, 65(2):257–268, 2023. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Ming:2023:DGP

[MWG23] Deyu Ming, Daniel Williamson, and Serge Guillas. Deep Gaussian process emulation using stochastic imputation. Technometrics, 65 (2):150–161, 2023. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Myers:2021:BRM

[Mye21] Donald E. Myers. Book review: Modelling Spatial and Spatial-Temporal Data: A Bayesian Approach, by Robert Haining and Guanquan Li. Boca Raton, FL: CRC Press, Taylor and Francis, 2020, xxvi + 597 pp., $119.35, ISBN: 13-4822-3742-9. Technometrics, 62(4):561, 2021. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Mak:2022:TFO

[MZHW22] Simon Mak, Yuanshuo Zhao, Lavonne Hoang, and C. F. Jeff Wu. TSEC: a framework for online experimentation under experimental constraints. Technometrics, 64(4):513–523, 2022. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Ndolo:2024:BRS

[Ndo24] Antony Ndolo. Book review: Smart Grid and Enabling Technologies, Shady S. Refaat, Omar Ellabban, Sertac Bayhan, Haitham Abu-Rub, Dr. Frede Blaabjerg, Miroslav M. Begovic, and Hoboken, NJ: Wiley-IEEE Press, 2021, 512 pp, $140.00, ISBN 978-1-119-42231-0. Technometrics, 66(1):137–139, 2024. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Nelson:2021:BRS

[Nel21] Wayne B. Nelson. Book review: Statistical Methods for Reliability Data, Second Edition. Edited by William Q. Meeker, Luís A. Escobar, and Francis G. Pascual, New York, Wiley, 2021, xiii + 545 pp., $135 (Hardback), ISBN: 978-1-118-11545-9, ePub ISBN 978-1-118-59448-3 for Kindle readers, ePDF ISBN 978-1-118-59459-9. Technometrics, 63(3):437–440, 2021. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

NunezAres:2020:EMS

[NG20] José Núñez Ares and Peter Goos. Enumeration and multicriteria selection of orthogonal minimally aliased response surface designs. Technometrics, 62(1):21–36, 2020. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Nguyen:2020:CEM

[NPM20] Nam-Ky Nguyen, Tung-Dinh Pham, and Phuong Vuong Mai. Constructing D-efficient mixed-level foldover designs using Hadamard matrices. Technometrics, 62(1):48–56, 2020. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Oates:2022:SAS

[OKF22] Chris J. Oates, Wilfrid S. Kendall, and Liam Fleming. A statistical approach to surface metrology for 3D-printed stainless steel. Technometrics, 64(3):370–383, 2022. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Olive:2020:BRS

[Oli20] David J. Olive. Book review: Statistical Programming in SAS (2nd ed.), by A. J. Bailer. Boca Raton, FL: Chapman & Hall/ CRC, Taylor & Francis Group, 2020, xv + 362 pp., $79.95, ISBN: 978-0-367-35797-9. Technometrics, 62(2):1–288, 2020. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Olive:2022:BRS

[Oli22] David J. Olive. Book review: Stochastic Processes with R: an Introduction by O. Korosteleva, Boca Raton, FL: Chapman and Hall, CRC Press, 2022, x + 190 pp., $99.95, ISBN 978-1-032-15373-5. Technometrics, 64(4):579, 2022. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Peng:2020:SPD

[PC20] Chien-Yu Peng and Ya-Shan Cheng. Student-t processes for degradation analysis. Technometrics, 62(2):223–235, 2020. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Pope:2021:GPM

[PGB^{+}21] Christopher A. Pope, John Paul Gosling, Stuart Barber, Jill S. Johnson, Takanobu Yamaguchi, Graham Feingold, and Paul G. Blackwell. Gaussian process modeling of heterogeneity and discontinuities using Voronoi tessellations. Technometrics, 63(1):53– 63, 2021. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Passino:2022:SCS

[PHRD22] Francesco Sanna Passino, Nicholas A. Heard, and Patrick Rubin-Delanchy. Spectral clustering on spherical coordinates under the degree-corrected stochastic blockmodel. Technometrics, 64(3): 346–357, 2022. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Pourmohamad:2021:SFA

[PL21] Tony Pourmohamad and Herbert K. H. Lee. The statistical filter approach to constrained optimization. Technometrics, 62(3):303– 312, 2021. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Pourmohamad:2021:SGP

[Pou21] Tony Pourmohamad. Surrogates: Gaussian process modeling, design, and optimization for the applied sciences. Technometrics, 63 (1):144–145, 2021. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Pratola:2022:ESI

[Pra22] Matthew T. Pratola. Editorial: Special issue on Industry 4.0. Technometrics, 64(4):435–436, 2022. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Prakash:2022:GPA

[PTD22] Abhinav Prakash, Rui Tuo, and Yu Ding. Gaussian process-aided function comparison using noisy scattered data. Technometrics, 64(1):92–102, 2022. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Prakash:2023:TOP

[PTD23] Abhinav Prakash, Rui Tuo, and Yu Ding. The temporal overfitting problem with applications in wind power curve modeling. Technometrics, 65(1):70–82, 2023. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Qu:2021:BDU

[QBN21] Kai Qu, Jonathan R. Bradley, and Xufeng Niu. Boundary detection using a Bayesian hierarchical model for multiscale spatial data. Technometrics, 63(1):64–76, 2021. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Qiu:2020:NPC

[QLL20] Peihua Qiu, Wendong Li, and Jun Li. A new process control chart for monitoring short-range serially correlated data. Technometrics, 62(1):71–83, 2020. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Qadir:2021:ESD

[QSK21] Ghulam A. Qadir, Ying Sun, and Sebastian Kurtek. Estimation of spatial deformation for nonstationary processes via variogram alignment. Technometrics, 63(4):548–561, 2021. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Qiu:2022:TSL

[QX22] Peihua Qiu and Xiulin Xie. Transparent sequential learning for statistical process control of serially correlated data. Technometrics, 64(4):487–501, 2022. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Qiu:2020:PMR

[QXY20] Peihua Qiu, Zhiming Xia, and Lu You. Process monitoring ROC curve for evaluating dynamic screening methods. Technometrics, 62(2):236–248, 2020. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Rahmouni:2023:BRD

[Rah23] Mohieddine Rahmouni. Book review: Dark Data: Why What You Don’t Know Matters by David J. Hand, Princeton, NJ: Princeton University Press, 2020, xii + 330 pp., ISBN 978-0-691-18237-7. Technometrics, 65(1):129–131, 2023. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Rougier:2024:BMI

[RD24] Jonathan Rougier and Andrew Duncan. Bayesian modeling and inference for one-shot experiments. Technometrics, 66(1):55–64, 2024. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Ryan:2023:DCC

[RK23] Sean Ryan and Rebecca Killick. Detecting changes in covariance via random matrix theory. Technometrics, 65(4):480–491, 2023. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Raymaekers:2021:FRC

[RR21] Jakob Raymaekers and Peter J. Rousseeuw. Fast robust correlation for high-dimensional data. Technometrics, 63(2):184–198, 2021. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Raymaekers:2022:CMV

[RRH22] Jakob Raymaekers, Peter J. Rousseeuw, and Mia Hubert. Class maps for visualizing classification results. Technometrics, 64(2): 151–165, 2022. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Saadah:2023:BRA

[Sa’23a] Aminatus Sa’adah. Book review: Artificial Intelligence with Python, Teik Toe Teoh and Zheng Rong, Springer, 2022, 336 pp., EUR 16.99, ISBN 978-981-16-8615-3. Technometrics, 65(3):451– 452, 2023. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Shi:2023:PFL

[SA23b] Naichen Shi and Raed Al Kontar. Personalized federated learning via domain adaptation with an application to distributed 3D printing. Technometrics, 65(3):328–339, 2023. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Suprianto:2023:BRM

[SA23c] Suprianto and Nur Alam. Book review: Mathematical Modeling in Biology: A Research Methods Approach, 1st ed. Shandelle M. Henson and James L. Hayward, Boca Raton, FL: Chapman and Hall/CRC Press, 2023, XXI + 315 pp., ISBN: 978-1-032-20821-3 (hbk). Scope: Textbook. Level: Advanced Mathematics and Science Undergraduate Students and Graduate-Level Biology Students. Technometrics, 65(3):448–449, 2023. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Stallrich:2023:OSD

[SAMJ23] Jonathan Stallrich, Katherine Allen-Moyer, and Bradley Jones. D- and A-optimal screening designs. Technometrics, 65(4):492– 501, 2023. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Sauter:2021:BRA

[Sau21] Roger M. Sauter. Book review: Advanced R (2nd ed.). Technometrics, 62(3):417, 2021. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Sauter:2023:BRF

[Sau23] Roger Sauter. Book review: Foundations of Statistics for Data Scientists: With R and Python by Alan Agresti and Maria Kateri, Boca Raton, FL: Chapman and Hall/CRC, 2022, xv + 453 pp., $136.46 (hcb). Technometrics, 65(1):132–133, 2023. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Stroh:2022:SDM

[SBD^{+}22] Rémi Stroh, Julien Bect, Séverine Demeyer, Nicolas Fischer, Damien Marquis, and Emmanuel Vazquez. Sequential design of multi-fidelity computer experiments: Maximizing the rate of stepwise uncertainty reduction. Technometrics, 64(2):199–209, 2022.

CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Striegel:2022:MFF

[SBWK22] Christoph Striegel, Jonas Biehler, Wolfgang A. Wall, and Göran Kauermann. A multifidelity function-on-function model applied to an abdominal aortic aneurysm. Technometrics, 64(3):279–290, 2022. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Sadeghi:2021:SDM

[SCA21] Soheil Sadeghi, Peter Chien, and Neeraj Arora. Sliced designs for multi-platform online experiments. Technometrics, 62(3):387– 402, 2021. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Seaver:2020:BRT

[Sea20] William L. Seaver. Book review: Time Series: A First Course With Bootstrap Starter, by Tucker S. McElroy and Dimitris N. Politis. Boca Raton, FL: CRC Press, Taylor & Francis Group, 2020, xvii + 566 pp., $99.95, ISBN: 978-1-4398-7651-0. Technometrics, 62(2):1–289, 2020. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Sadeghi:2022:NPL

[SGG^{+}22] Soheil Sadeghi, Somit Gupta, Stefan Gramatovici, Jiannan Lu, Hao Ai, and Ruhan Zhang. Novelty and primacy: a long-term estimator for online experiments. Technometrics, 64(4):524–534, 2022. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Sauer:2023:ALD

[SGH23] Annie Sauer, Robert B. Gramacy, and David Higdon. Active learning for deep Gaussian process surrogates. Technometrics, 65(1):4–18, 2023. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Shen:2020:AHM

[SKD20] Sumin Shen, Lulu Kang, and Xinwei Deng. Additive heredity model for the analysis of mixture-of-mixtures experiments. Technometrics, 62(2):265–276, 2020. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Singh:2023:STL

[SS23] Rakhi Singh and John Stufken. Selection of two-level supersaturated designs for main effects models. Technometrics, 65(1):96– 104, 2023. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Scimone:2022:SMM

[STC^{+}22] Riccardo Scimone, Tommaso Taormina, Bianca Maria Colosimo, Marco Grasso, Alessandra Menafoglio, and Piercesare Secchi. Statistical modeling and monitoring of geometrical deviations in complex shapes with application to additive manufacturing. Technometrics, 64(4):437–456, 2022. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Shen:2022:SRM

[SWL^{+}22] Bo Shen, Rongxuan Wang, Andrew Chung Chee Law, Rakesh Kamath, Hahn Choo, and Zhenyu (James) Kong. Super resolution for multi-sources image stream data using smooth and sparse tensor completion and its applications in data acquisition of additive manufacturing. Technometrics, 64(1):2–17, 2022. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Sun:2021:SMM

[SYH21] Qiuzhuang Sun, Zhi-Sheng Ye, and Yili Hong. Statistical modeling of multivariate destructive degradation tests with blocking. Technometrics, 62(4):536–548, 2021. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Thomas:2023:FDH

[TCXM23] Andrew M. Thomas, Peter A. Crozier, Yuchen Xu, and David S. Matteson. Feature detection and hypothesis testing for extremely noisy nanoparticle images using topological data analysis. Technometrics, 65(4):590–603, 2023. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Tsai:2023:DOA

[Tsa23] Shin-Fu Tsai. Dual-orthogonal arrays for order-of-addition twolevel factorial experiments. Technometrics, 65(3):388–395, 2023. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Vorechovsky:2021:MMP

[VE21] Miroslav Vorechovský and Jan Eliás. Modification of the maximin and φ_{p} (phi) criteria to achieve statistically uniform distribution of sampling points. Technometrics, 62(3):371–386, 2021. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Vazquez:2020:PDS

[VGS20] Alan R. Vazquez, Peter Goos, and Eric D. Schoen. Projections of definitive screening designs by dropping columns: Selection and evaluation. Technometrics, 62(1):37–47, 2020. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Vishwakarma:2021:BRS

[VL21a] Srishti Vishwakarma and Vyacheslav Lyubchich. Book review: Spatial Ecology and Conservation Modeling: Applications With R. Technometrics, 62(3):416–417, 2021. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Vishwakarma:2021:BRT

[VL21b] Srishti Vishwakarma and Vyacheslav Lyubchich. Book review: Time Series Clustering and Classification, by Elizabeth Ann Maharaj, Pierpaolo D’Urso, and Jorge Caiado, Boca Raton, FL: Chapman & Hall/CRC Press, 2019, xv + 228 pp., $174.95 (Hardback), ISBN: 978-1-49-877321-8. Technometrics, 63(3):441, 2021. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Wycoff:2022:SPL

[WBG22] Nathan Wycoff, Mickaël Binois, and Robert B. Gramacy. Sensitivity prewarping for local surrogate modeling. Technometrics, 64 (4):535–547, 2022. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Wu:2021:CVC

[WDWZ21] Qian Wu, Xinwei Deng, Shiren Wang, and Li Zeng. Constrained varying-coefficient model for time-course experiments in soft tissue fabrication. Technometrics, 63(2):249–262, 2021. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Wang:2022:RFP

[WDZS22] Andi Wang, Juan Du, Xi Zhang, and Jianjun Shi. Ranking features to promote diversity: an approach based on sparse distance correlation. Technometrics, 64(3):384–395, 2022. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Wilson:2021:ASS

[WF21] Kevin J. Wilson and Malcolm Farrow. Assurance for sample size determination in reliability demonstration testing. Technometrics, 63(4):523–535, 2021. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Wang:2024:SDF

[WGKJ24] Shangkun Wang, Adam P. Generale, Surya R. Kalidindi, and V. Roshan Joseph. Sequential designs for filling output spaces. Technometrics, 66(1):65–76, 2024. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Wludyka:2021:BRA

[Wlu21] Peter Wludyka. Book review: Artificial Intelligence for Drug Development, Precision Medicine, and Healthcare, by Mark Chang. Boca Raton, FL: CRC Taylor and Francis, 2020, xii + 345 pp., $107.78 (hardback), $46.36 (eBook), ISBN: 978-0-367-36292-8. Technometrics, 62(4):561–562, 2021. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Wludyka:2023:BRP

[Wlu23] Peter Wludyka. Book review: Principles of Biostatistics, by Marcello Pagano, Kimberlee Gauvreau, and Heather Mattie, CRC Press, Boca Raton, FL, 2022, 620 pp., $79.95, ISBN 978-0-429-34051-2. Technometrics, 65(2):298–299, 2023. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Weaver:2021:BMP

[WM21] Brian P. Weaver and William Q. Meeker. Bayesian methods for planning accelerated repeated measures degradation tests. Technometrics, 63(1):90–99, 2021. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Weese:2021:SSS

[WSSE21] Maria L. Weese, Jonathan W. Stallrich, Byran J. Smucker, and David J. Edwards. Strategies for supersaturated screening: Group orthogonal and constrained Var(s) designs. Technometrics, 63(4): 443–455, 2021. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Winkel:2021:SOL

[WSSR21] Munir A. Winkel, Jonathan W. Stallrich, Curtis B. Storlie, and Brian J. Reich. Sequential optimization in locally important dimensions. Technometrics, 63(2):236–248, 2021. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Weiss:2023:NCC

[WT23] Christian H. Weiß and Murat Caner Testik. Nonparametric control charts for monitoring serial dependence based on ordinal patterns. Technometrics, 65(3):340–350, 2023. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Wu:2021:CRR

[Wu21] Yichao Wu. Can’t ridge regression perform variable selection? Technometrics, 63(2):263–271, 2021. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Wang:2021:SPC

[WV21] Yixin Wang and Stefan Van Aelst. Sparse principal component analysis based on least trimmed squares. Technometrics, 62(4): 473–485, 2021. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Weaver:2023:AET

[WV23] Brian P. Weaver and Scott A. Vander Wiel. Accelerated event times with multiple thresholds. Technometrics, 65(4):502–513, 2023. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Wu:2024:STA

[WWYC24] Tong Wu, Yudong Wang, Zhi-Sheng Ye, and Nan Chen. Spatiotemporal analysis and prediction of mass telecommunication base station failure events. Technometrics, 66(1):77–89, 2024. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Wu:2021:AMC

[WXJT21] Jianguo Wu, Honglun Xu, Feng Ju, and Tzu-Liang (Bill) Tseng. Adaptive minimum confidence region rule for multivariate initialization bias truncation in discrete-event simulations. Technometrics, 62(4):499–512, 2021. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Wang:2023:SCP

[WXX^{+}23] Haoyun Wang, Liyan Xie, Yao Xie, Alex Cuozzo, and Simon Mak. Sequential change-point detection for mutually exciting point processes. Technometrics, 65(1):44–56, 2023. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Wang:2021:AIM

[WY21] Qin Wang and Xiangrong Yin. Aggregate inverse mean estimation for sufficient dimension reduction. Technometrics, 63(4):456– 465, 2021. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Wang:2022:FEL

[WZ22] Boxiang Wang and Hui Zou. Fast and exact leave-one-out analysis of large-margin classifiers. Technometrics, 64(3):291–298, 2022. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Wu:2022:MHM

[WZL22] Hui Wu, Chen Zhang, and Yan-Fu Li. Monitoring heterogeneous multivariate profiles based on heterogeneous graphical model. Technometrics, 64(2):210–223, 2022. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Wang:2022:DDD

[WZQ22] Guanghui Wang, Changliang Zou, and Peihua Qiu. Data-driven determination of the number of jumps in regression curves. Technometrics, 64(3):312–322, 2022. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Wang:2023:CCR

[WZZ23] Xin Wang, Xin Zhang, and Zhengyuan Zhu. Clustered coefficient regression models for Poisson process with an application to seasonal warranty claim data. Technometrics, 65(4):514–523, 2023. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Xiong:2021:RAI

[Xio21] Shifeng Xiong. The reconstruction approach: From interpolation to regression. Technometrics, 63(2):225–235, 2021. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Xiao:2023:MOF

[XJR23] Qian Xiao, V. Roshan Joseph, and Douglas M. Ray. Maximum one-factor-at-a-time designs for screening in computer experiments. Technometrics, 65(2):220–230, 2023. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Xiao:2021:NBM

[XKSK21] Sai Xiao, Athanasios Kottas, Bruno Sansó, and Hyotae Kim. Nonparametric Bayesian modeling and estimation for renewal processes. Technometrics, 63(1):100–115, 2021. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Xie:2024:GFR

[XQ24] Xiulin Xie and Peihua Qiu. A general framework for robust monitoring of multivariate correlated processes. Technometrics, 66(1): 40–54, 2024. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Xiang:2022:RPS

[XQWL22] Dongdong Xiang, Peihua Qiu, Dezhi Wang, and Wendong Li. Reliable post-signal fault diagnosis for correlated high-dimensional data streams. Technometrics, 64(3):323–334, 2022. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Xu:2023:JDH

[XSL^{+}23] Feng Xu, Lianjie Shu, Yanting Li, Binhui, and Wang. Joint diagnosis of high-dimensional process mean and covariance matrix based on Bayesian model selection. Technometrics, 65(4):465– 479, 2023. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Xu:2023:OSC

[XWYL23] Ruiyu Xu, Jianguo Wu, Xiaowei Yue, and Yongxiang Li. Online structural change-point detection of high-dimensional streaming data via dynamic sparse subspace learning. Technometrics, 65(1): 19–32, 2023. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Xian:2021:SMR

[XYWL21] Xiaochen Xian, Honghan Ye, Xin Wang, and Kaibo Liu. Spatiotemporal modeling and real-time prediction of origindestination traffic demand. Technometrics, 63(1):77–89, 2021.

CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Yue:2021:JME

[YA21] Xubo Yue and Raed Al Kontar. Joint models for event prediction from time series and survival data. Technometrics, 63(4):477– 486, 2021. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Yang:2024:DOT

[Yan24] Jiannan Yang. Decision-oriented two-parameter Fisher information sensitivity using symplectic decomposition. Technometrics, 66(1):28–39, 2024. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Ye:2024:SME

[YCTY24] Xin Ye, Jiaxiang Cai, Loon Ching Tang, and Zhi-Sheng Ye. Statistical modeling of the effectiveness of preventive maintenance for repairable systems. Technometrics, 66(1):118–130, 2024. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Yu:2022:HDC

[YFL22] Guan Yu, Haoda Fu, and Yufeng Liu. High-dimensional costconstrained regression via nonconvex optimization. Technometrics, 64(1):52–64, 2022. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Yoon:2023:LNR

[YK23] Heegeon Yoon and Heeyoung Kim. Label-noise robust deep generative model for semi-supervised learning. Technometrics, 65(1): 83–95, 2023. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Yue:2020:TME

[YPLS20] Xiaowei Yue, Jin Gyu Park, Zhiyong Liang, and Jianjun Shi. Tensor mixed effects model with application to nanomanufacturing inspection. Technometrics, 62(1):116–129, 2020. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

You:2020:EMO

[YQ20] Lu You and Peihua Qiu. An effective method for online disease risk monitoring. Technometrics, 62(2):249–264, 2020. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Yang:2021:APM

[YQ21] Kai Yang and Peihua Qiu. Adaptive process monitoring using covariate information. Technometrics, 63(3):313–328, 2021. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Yin:2023:GMF

[YSM23] Hang Yin, Abolfazl Safikhani, and George Michailidis. A general modeling framework for network autoregressive processes. Technometrics, 65(4):579–589, 2023. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Yang:2021:CPM

[YSX21] Jian-Feng Yang, Fasheng Sun, and Hongquan Xu. A componentposition model, analysis and design for order-of-addition experiments. Technometrics, 63(2):212–224, 2021. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Yuan:2021:CRR

[Yua21] Ming Yuan. Comment: From ridge regression to methods of regularization. Technometrics, 62(4):447–450, 2021. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Yao:2023:ASM

[YXW23] Jinwei Yao, Xiaochen Xian, and Chao Wang. Adaptive sampling for monitoring multi-profile data with within-and-between profile correlation. Technometrics, 65(3):375–387, 2023. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Zhang:2023:CDM

[ZBA23] Kungang Zhang, Anh T. Bui, and Daniel W. Apley. Concept drift monitoring and diagnostics of supervised learning models via score vectors. Technometrics, 65(2):137–149, 2023. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Zhang:2021:DDD

[ZCG21] Boya Zhang, D. Austin Cole, and Robert B. Gramacy. Distancedistributed design for Gaussian process surrogates. Technometrics, 63(1):40–52, 2021. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Zhao:2021:IGA

[ZdC21] Xueqi Zhao and Enrique del Castillo. An intrinsic geometrical approach for statistical process control of surface and manifold data. Technometrics, 63(3):295–312, 2021. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Zelterman:2022:BRH

[Zel22] Daniel Zelterman. Book review: Handbook of Statistical Methods for Randomized Controlled Clinical Trials by Kyung Mann Kim, Frand Bretz, Ying Kuen Cheung, and Lisa V Hampson, eds., Boca Raton, FL, CRC/Chapman & Hall, 2022, 620 pp., $185.15, ISBN 978-1-4987-1462-4. Technometrics, 64(3):424–425, 2022. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Zulfaidil:2024:BRS

[ZGAR24] Zulfaidil, Utari Akhir Gusti, Waliyyatu Azzahra, and Gantina Rachmaputri. Book review: The Science of Hockey: The Math, Technology, and Data Behind the Sport, Edited by Kevin Snow, Foreword by John Vogl, New York, NY: Sports Publishing, 2023, 224 pp., $19.99 (ppb), ISBN 978-1-68358-466-7. Technometrics, 66(1):136–137, 2024. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Zhang:2021:SMB

[ZH21] Qiong Zhang and Youngdeok Hwang. Sequential model-based optimization for continuous inputs with finite decision space. Technometrics, 62(4):486–498, 2021. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Zhao:2022:APO

[ZHMY22] Xinyu Zhao, Jiuyun Hu, Yajun Mei, and Hao Yan. Adaptive partially observed sequential change detection and isolation. Technometrics, 64(4):502–512, 2022. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Zhang:2022:LOD

[ZK22] Qiong Zhang and Lulu Kang. Locally optimal design for A/B tests in the presence of covariates and network dependence. Technometrics, 64(3):358–369, 2022. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Zhang:2023:BCP

[ZM23] Wanrong Zhang and Yajun Mei. Bandit change-point detection for real-time monitoring high-dimensional data under sampling control. Technometrics, 65(1):33–43, 2023. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Zheng:2023:PNO

[ZMXX23] Xiaojun Zheng, Simon Mak, Liyan Xie, and Yao Xie. PER-CEPT: A new online change-point detection method using topological data analysis. Technometrics, 65(2):162–178, 2023. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Zou:2021:CRR

[Zou21] Hui Zou. Comment: Ridge regression — still inspiring after 50 years. Technometrics, 62(4):456–458, 2021. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Zhen:2023:IBF

[ZPS23] Zhong Zhen, Kamran Paynabar, and Jianjun (Jan) Shi. Imagebased feedback control using tensor analysis. Technometrics, 65 (3):305–314, 2023. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Zhang:2024:FRL

[ZSD24] Maoyu Zhang, Yan Song, and Wenlin Dai. Fast robust location and scatter estimation: a depth-based method. Technometrics, 66(1):14–27, 2024. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Zhu:2021:OCR

[ZSPT21] Xiaochen Zhu, Martin Slawski, P. Jonathon Phillips, and Liansheng Larry Tang. Order-constrained ROC regression with application to facial recognition. Technometrics, 63(3):343–353, 2021. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Zhang:2021:LVA

[ZTCA21] Yichi Zhang, Siyu Tao, Wei Chen, and Daniel W. Apley. A latent variable approach to Gaussian process modeling with qualitative and quantitative factors. Technometrics, 62(3):291–302, 2021. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Zan:2023:SRB

[ZWX23] Xin Zan, Di Wang, and Xiaochen Xian. Spatial rank-based augmentation for nonparametric online monitoring and adaptive sampling of big data streams. Technometrics, 65(2):243–256, 2023. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Zhai:2023:MSD

[ZY23] Qingqing Zhai and Zhi-Sheng Ye. A multivariate stochastic degradation model for dependent performance characteristics. Technometrics, 65(3):315–327, 2023. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Zhang:2021:DMF

[ZYLS21] Chen Zhang, Hao Yan, Seungho Lee, and Jianjun Shi. Dynamic multivariate functional data modeling via sparse subspace learning. Technometrics, 63(3):370–383, 2021. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).